### v1

# Lyndon pairs and the lexicographically greatest perfect necklace

Abstract.
Fix a finite alphabet. A necklace is a circular word. For positive integers $n$ and $k$, a necklace is $(n,k)$-perfect if all words of length $n$ occur $k$ times but at positions with different congruence modulo $k$, for any convention of the starting position. We define the notion of a Lyndon pair and we use it to construct the lexicographically greatest $(n,k)$-perfect necklace, for any $n$ and $k$ such that $n$ divides $k$ or $k$ divides $n$. Our construction generalizes Fredricksen and Maiorana’s construction of the lexicographically greatest de Bruijn sequence of order $n$, based on the concatenation of the Lyndon words whose length divide $n$.
Key words and phrases:
de Bruijn sequences, Lyndon words
2020 Mathematics Subject Classification:
Primary 68R15, 05A05; Secondary 11K16
1. Introduction
Let $\Sigma$ be a finite alphabet with at least two symbols. A word on $\Sigma$ is a finite sequence of symbols, and a necklace is the equivalence class of a word under rotations. Given two positive integers, $n$ and $k$, a necklace is $(n,k)$-perfect if all words of length $n$ occur $k$ times but at positions with different congruence modulo $k$, for any convention of the starting position. The well known circular de Bruijn sequences of order $n$, see [3, 7, 8], that we call de Bruijn necklaces of order $n$, are exactly the $(n,k)$-perfect necklaces for $k=1$. For example, 1110010011100100 1110010011100100 is a $(2,2)$-perfect for $\Sigma=\{0,1\}$. The $(n,k)$-perfect necklaces correspond to Hamiltonian cycles in the tensor product of the de Bruijn graph with a simple cycle of length $k$.
A thorough presentation of perfect necklaces appears in [1]. With the purpose of constructing normal numbers with very fast convergence to normality M. Levin in [9] gives two constructions of perfect necklaces. One based on arithmetic progressions with difference coprime with the alphabet size which yields a $(n,n)$-perfect necklaces. The other based on Pascal triangle matrix which yields nested $(n,n)$-perfect necklaces when $n$ is a power of 22 22. In [2] there is a method of constructing all nested $(n,n)$-perfect necklaces.
Assume the lexicographic order on words. A Lyndon word is a nonempty aperiodic word that is lexicographically greater than all of its rotations. For example, the Lyndon words over alphabet $\{0,1\}$ sorted by length and then in decreasing lexicographical order within each length are
1,0,10,110,100,1110,1100,1000,11110,11100,11010,11000,10100,10000,1010110100111011001000111101110011010110001010010000 1,$0$,...1 , 0 , 10 , 110 , 100 , 1110 , 1100 , 1000 , 11110 , 11100 , 11010 , 11000 , 10100 , 10000 , …
Lyndon words were introduced by Lyndon in mid 1950s [10, 11]. They provide a nice factorization of the free monoid: each word $w$ of the free monoid $\Sigma^{*}$ has a unique decomposition as a product $w=u_{1}\ldots u_{n}$ of a non-increasing sequence of Lyndon words $u_{1},\ldots u_{n}$ in the lexicographic order. The problem to compute the prime factorization of a given word has a solution in time linear to the length of the given word [4], see also [6]. Fredricksen and Maiorana [5] showed that the concatenation in decreasing lexicographic order of all the Lyndon words whose length divides a given positive integer number $n$, yields a de Bruijn necklace of order $n$. For example, the concatenation of the binary Lyndon words whose length divides four is (the spaces are for ease of reading),
1 1110 1100 10 1000 0.1111011001010000 1\ 1110\ 1100\ 10\ 1000\ 0.1 1110 1100 10 1000 0 .
This construction, together with the efficient generation of Lyndon words, provides an efficient method for constructing the lexicographically greatest de Bruijn necklace of each order $n$ in linear time and logarithmic space.
In this note we present the notion of Lyndon pairs and we use it to generalize Fredricksen and Maiorana’s algorithm to construct the lexicographically greatest $(n,k)$-perfect necklaces, for any $n$ and $k$ such that $n$ divides $k$ or $k$ divides $n$.
2. Lyndon pairs in lexicographical order
2.1. Lyndon pairs
We assume a finite alphabet $\Sigma$ with cardinality $s$, with $s\geq 2$. Without loss of generality we use $\Sigma=\{0,\ldots,s-1\}$. We use lowercase letters $a,b,c$ possibly with subindices for alphabet symbols. Words are finite sequences of symbols that we write $a_{1}a_{2}..a_{n}$ or with a capital letter $A,B,C$. We write $a^{\ell}$ to denote the word of length $\ell$ made just of $a^{\prime}$s. and we write $A^{\ell}$ to denote the word made of $\ell$ copies of $A$. The concatenation of two words $A$ and $B$ is written $AB$. The length of a word $A$ is denoted with $|A|$. The positions of a word $A$ are numbered from 11 11 to $|A|$. We use $>$ to denote the decreasing lexicographic order on words and we write $A\geq B$ with the when $A>B$ or $A=B$.
We use lowercase letters $h,\ldots,z$ to denote non-negative integers. We write $k|n$ to say that $k$ divides $n$. We write $\mathbb{Z}_{k}$ for the set $\mathbb{Z}/k\mathbb{Z}$ of residues modulo $k$. We also use $<$ and $>$ for the natural orders on $\mathbb{Z}_{k}$ and ${\mathcal{N}}$ and, as usual, $u\leq v$ when $u<v$ or $u=v$; and $v\geq u$ when $v>u$ or $v=u$. When $u<v$ we may write $v>u$
We work with pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$ when $k|n$ or $n|k$. This condition on $n$ and $k$ is assumed all along the sequel. We refer to pairs $\langle A,u\rangle$ with calligraphic letter ${\mathcal{A}},{\mathcal{B}},{\mathcal{C}},\ldots$ We consider the following order $\succ$ over $\Sigma^{n}\times\mathbb{Z}_{k}$.
Definition (order $\succ$ on $\Sigma^{n}\times\mathbb{Z}_{k}$).
$\langle A,u\rangle\succ\langle B,v\rangle\text{ exactly when either }(A>B\text{ and }u=v)\text{ or }(k-1-u)>(k-1-v).$
The smallest the second component in $\mathbb{Z}_{k}$, the $\succ$-greater the pair. Among pairs with the same second component in $\mathbb{Z}_{k}$, the order $\succ$ is defined with the decreasing lexicographic order on $\Sigma^{n}$. Thus, $\langle(s-1)^{n},0\rangle$ is the $\succ$-greatest in $\Sigma^{n}\times\mathbb{Z}_{k}$. As usual, we write $\langle A,u\rangle\succeq\langle B,v\rangle$ exactly when $\langle A,u\rangle\succ\langle B,v\rangle$ or $\langle A,u\rangle=\langle B,v\rangle$.
It is possible to concatenate pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$ having the same second component: the concatenation of $\langle A,u\rangle$ and $\langle B,u\rangle$ is $\langle AB,u\rangle$. We write $\langle A,u\rangle^{t}$ to denote the pair $\langle A^{t},tu\rangle$.
Definition (rotation of a pair).
Given a pair $\langle a_{1}..a_{n},u\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ its left rotation is the pair $\langle a_{n}a_{1}..a_{n-1},u-1\rangle$ and its right rotation is the pair $\langle a_{2}..a_{n}a_{1},u+1\rangle$.
For $s=3$, $n=k=5$ the right rotation of pair $\langle 13212,4\rangle$, is $\langle 21321,0\rangle$, and its left rotation is $\langle 32121,3\rangle$. The rotation function induces a relation between pairs: two pairs are related if successive rotations initially applied to the first yield the second. This relation is clearly reflexive and transitive. For pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$, when $k|n$ or $n|k$, the rotation has an inverse, given by successive rotations. So the relation is also symmetric, hence, an equivalence relation.
Definition (maximal pair).
A necklace in $\Sigma^{n}\times\mathbb{Z}_{k}$ is a set of pairs that are equivalent under rotations,
$\{\langle a_{i+1}..a_{n}a_{1}...a_{i},u+i\rangle:0\leq i<\max(n,k)\}$
In each necklace we are interested in the pair that is maximal in the order $\succ$. We call it maximal .
For example, for $n=4$ and $k=2$ the pair $\langle 1110,0\rangle$ is maximal among its rotations, hence $\langle 1011,0\rangle$ is not maximal and $\langle 1101,1\rangle$ is not maximal either. The pair $\langle 1010,0\rangle$ is maximal among its rotations.
Observation 1.
When $n|k$ all pairs $\langle A,0\rangle$ for $A\in\Sigma^{n}$ are maximal .
Observation 2.
Let $n$ and $k$ positive integers such that $k|n$ or $n|k$. Then, ${\mathcal{A}}\in\Sigma^{n}\times\mathbb{Z}_{k}$ is maximal exactly when ${\mathcal{A}}^{p}$, for any $p\geq 1$, is maximal .
Proof.
($\implies$). In case $n|k$ all pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$ with second component 00 0 are maximal . We prove it for $k|n$. Assume ${\mathcal{A}}^{p}$ is maximal rotation, but ${\mathcal{A}}$ is not. Then ${\mathcal{A}}$ has a rotation
${\mathcal{R}}=\langle({a}_{j+1}\cdots{a}_{n}{a}_{1}\cdots{a}_{j}),0\rangle$
where $k|j$ such that ${\mathcal{R}}\succ{\mathcal{A}}$. But this implies that
$\langle{a}_{j+1}\cdots{a}_{n}{a}_{1}\cdots{a}_{j},0\rangle^{p}\succ\langle{a}_{1}\cdots{a}_{n},0\rangle^{p},$
contradicting that ${\mathcal{A}}^{p}$ is maximal .
[…]
Observation 3.
If ${\mathcal{A}}$ is a maximal pair then none of its rotations are $\succ$-greater than ${\mathcal{A}}$, but there may be a rotation that is equal to ${\mathcal{A}}$.
For a word $A=a_{1}..a_{n}$ we write $A_{i}$ to denote its prefix of length $i$, that is, $a_{1}..a_{i}$.
Lemma 1.
Let $n$ and $k$ be positive integers such that $k|n$ or $n|k$. Let ${\mathcal{A}}=\langle A,0\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ be maximal and different from $\langle 0^{n},0\rangle$. Suppose $A=a_{1}...a_{n}$, let $i$ be such that $a_{i}>0$ and let
${\mathcal{B}}=\langle A_{i-1}(a_{i}-1)(s-1)^{j-i},0\rangle,$
where if $n|k$ then $j=n$; and if $k|n$ then $j$ is the smallest such that $k|j$ and $i\leq j<n$. Then, ${\mathcal{B}}$ is maximal .
Proof.
Assume ${\mathcal{A}}$ is a maximal pair and, by way of contradiction, assume that ${\mathcal{B}}=\langle A_{i-1}(a_{i}-1)(s-1)^{j-i},0\rangle$ is not a maximal pair. Then there is some $\ell$ multiple of $k$ such that :
$\langle{a}_{\ell+1}\cdots{a}_{i-1}(a_{i}-1)(s-1)^{j-i}A_{\ell},0+\ell\rangle\succ\langle A_{i-1}(a_{i}-1)(s-1)^{j-i},0\rangle$
Necessarily
${a}_{\ell+1}\cdots{a}_{i-1}\geq{a}_{1}\cdots{a}_{(i-1)-(\ell+1)+1}$
Since ${\mathcal{A}}$ is a maximal pair,
${a}_{1}\cdots{a}_{(i-1)-(\ell+1)+1}\geq{a}_{\ell+1}\cdots{a}_{i-1}.$
Therefore, $a_{\ell+1}=a_{1}$, $a_{\ell+2}=a_{2}$, $\dots$, $a_{i-1}=a_{(i-1)-(\ell+1)+1}$. This implies, $a_{i}-1\geq a_{(i-1)-(\ell+1)+2}$. Consequently,
${a}_{\ell+1}\cdots{a}_{i-1}a_{i}>{a}_{1}\cdots{a}_{(i-1)-(\ell+1)+2},$
but this contradicts that ${\mathcal{A}}$ es a maximal pair. ∎
For example, let $s=7$, $n=6$, $k=3$. Since ${\mathcal{A}}=\langle 456123,0\rangle$ is maximal , then ${\mathcal{B}}=\langle 455,0\rangle$ is maximal , defined by taking $i=3$, which is multiple of $k$, and there is no filling with $s-1=6$. Also ${\mathcal{B}}=\langle 366,0\rangle$ is maximal , because we take $i=1$ and we complete with $s-1=6$ up to position $j=3$.
We define the operator $\theta$ that given a pair in $\Sigma^{n}\times\mathbb{Z}_{k}$ with second component 00 0, but different from $\langle 0^{n},0\rangle$, it defines another pair in $\Sigma^{n}\times\mathbb{Z}_{k}$ with second component 00 0.
Definition (reduction).
Let $n$ and $k$ be positive integers such that $k|n$. The reduction of a word $A={a}_{1}\cdots{a}_{n}$, denoted $\overline{A}$, is the word ${a}_{1}\cdots{a}_{p}$ where $p$ is the smallest such that $k\mid p$, $p\mid n$ and ${a}_{1}\cdots{a}_{n}=({a}_{1}\cdots{a}_{p})^{n/p}$. The reduction of a pair ${\mathcal{A}}=\langle A,u\rangle$, denoted with $\overline{{\mathcal{A}}}$, is the pair $\langle\overline{A},u\rangle$.
When $k|n$, the reduction always exists because one can take $p=n$. Notice that if a pair in $\Sigma^{n}\times\mathbb{Z}_{k}$ is non reduced with period $p$, such that $k|p$ and $p|n$ then it has $n/p$ equal rotations. However, all the rotations of a reduced pair are pairwise different. For example, for $s=8$, $n=8$ and $k=2$, $\overline{\langle 10101010,0\rangle}=\langle 10,0\rangle$; $\overline{\langle 01230123,0\rangle}=\langle 0123,0\rangle$; $\overline{\langle 01234567,0\rangle}=\langle 01234567,0\rangle$.
Definition (expansion).
Let $n$ and $k$ be positive integers such that $n|k$. For $A\in\Sigma^{n}$, $\tilde{A}=A^{k/n}\in\Sigma^{k}$. For ${\mathcal{A}}=\langle A,u\rangle\in\Sigma^{n}\times\mathbb{Z}_{k}$, $\tilde{{\mathcal{A}}}=\langle\tilde{A},u\rangle\in\Sigma^{k}\times\mathbb{Z}_{k}$.
When $n|k$ the expansion always exists. For example for $s=3$, $n=2$ and $k=8$, $\tilde{\langle 12,0\rangle}=\langle 12121212,0\rangle$.
Definition (Lyndon pair).
Let $n$ and $k$ be positive integers. When $k|n$, the Lyndon pairs are the reductions of the maximal pairs in $\Sigma^{n}\times{\mathbb{Z}_{k}}$. When $n|k$ with $n<k$, the Lyndon pairs are the expansions of the maximal pairs $\langle A,0\rangle\in\Sigma^{n}\times\mathbb{Z}_{k}$.
Thus, when $k|n$, the Lyndon pairs are elements in $Sigma^{\leq n}\times\mathbb{Z}_{k}$. But when $n|k$ the Lyndon pairs are elements in $\Sigma^{n}\times\mathbb{Z}_{k}$.
Observation 4.
Each Lyndon pair is strictly $\succ$-greater than all of its rotations.
2.2. The operator $\theta$
Definition (operator $\theta$).
Let $n$ and $k$ be positive integers such that $k|n$ or $n|k$. For ${\mathcal{A}}=\langle A,0\rangle=\langle{a}_{1}\cdots{a}_{n},0\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ such that $a_{i}>a_{i+1}=\dots=a_{n}=0$, we define the operator $\theta:\Sigma^{n}\times\mathbb{Z}_{k}\to\Sigma^{n}\times\mathbb{Z}_{k}$,
$\theta\langle A,0\rangle=\langle[A_{i-1}(a_{i}-1)(s-1)^{j-i}]^{q}A_{n-qj},0\rangle,$
where,
if $n|k$ then $j=n$ and $q=1$; so, $\theta\langle A,0\rangle=\langle[A_{i-1}(a_{i}-1)(s-1)^{n-i},0\rangle$;
If $k|n$ then $j$ is the smallest integer such that $k|j$ and $j\geq i$, and $q$ is the greatest integer such that $q\leq n/j$. Thus, in either case, $j=i+((n-i)\mod k)$.
The operator $\theta$ is a function from $\Sigma^{n}\times\mathbb{Z}_{k}$ to $\Sigma^{n}\times\mathbb{Z}_{k}$. It is applicable on any pair with second component 00 0, except for $\langle 0^{n},0\rangle$. For example, for $s=2$, $n=6$ y $k=2$, $\theta\langle 010000,0\rangle=\langle 000000,0\rangle$; $\theta\langle 011000,0\rangle=\langle 010101,0\rangle$; $\theta\langle 011101,0\rangle=\langle 011100,0\rangle$.
Definition ($T$ yields the last).
Let $T$ be the integer such that $\theta^{T}\langle(s-1)^{n},0\rangle=\langle(0^{n},0\rangle)$.
Lemma 2.
The list $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$ is strictly decreasing in $\succ$.
Proof.
Let’s see that for every pair ${\mathcal{A}}=\langle A,0\rangle$, ${\mathcal{A}}\succ\theta{\mathcal{A}}$. Using the definition of $\theta$,
$\langle A,0\rangle\succ\langle[A_{i-1}(a_{i}-1)(s-1)^{j-i}]^{q}A_{n-qj},0\rangle.$
Since both pairs have second component 00 0, there is some $i$ such that
${a}_{1}\cdots{a}_{i}=A_{i-1}a_{i}>A_{i-1}(a_{i}-1).$
We conclude that $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$ is strictly decreasing in $\succ$. ∎
Observation 5.
When $n|k$ the operator $\theta$ yields a bijection between maximal pairs.
Proof.
Every pair of the form $\langle A,0\rangle$ is maximal because, when $k$ is a multiple of $n$, there is just this unique rotation. The definition of $\theta$ ensures that $\theta$ goes through all the pairs of the form $\langle A,0\rangle$. in $\succ$-decreasing order. The operator $\theta$ can be used forward for every pair $\langle A,0\rangle$ except for $\langle 0^{n},0\rangle$, and it can be used backwards for every pair $\langle A,0\rangle$ except for $\langle 0^{n},0\rangle$. Thus, except for the extremes, we can obtain the successor and the predecessor of a maximal pair in the order $\succ$. ∎
When $k|n$ the operator $\theta$ is not injective nor surjective over pairs with second component 00 0. For example, for $s=2$, $n=4$ and $k=2$, we see $\theta$ is not injective because $\theta\langle 0100,0\rangle=\langle 0000,0\rangle$ and also $\theta\langle 0001,0\rangle=\langle 0000,0\rangle$. To see that $\theta$ is not surjective observe that the pair ${\mathcal{B}}=\langle 1011,0\rangle$ is not in the image of $\theta$, because there is no pair ${\mathcal{A}}$ such that $\theta{\mathcal{A}}={\mathcal{B}}$.
It is possible to construct the reverse of list $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$. In case $n|k$ divides $k$ the operator $\theta$ yields a bijection between maximal pairs. In case $k|n$ and $k<n$, $\theta$ is not injective, there are pairs that have more than one preimage by $\theta$. However, except for $\langle(s-1)^{n},0\rangle$ every element has one predecessor in the list $(\theta^{i}(\langle s-1)^{n},0\rangle)_{0\leq i\leq T}$, which is just one of the possible preimages by $\theta$.
Lemma 3.
Let $n$ and $k$ positive integers such that $k|n$. Every element ${\mathcal{A}}=\langle A,0\rangle$ in $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$, except $\langle(s-1)^{n},0\rangle$, has a predecessor which is the preimage of ${\mathcal{A}}$ by $\theta$ given by
$\langle A_{u-1}(a_{u}+1))0^{n-u},0\rangle$
where
$A=\left(A_{u}(s-1)^{r-u}\right)^{w}A_{v}$
$a_{u}<(s-1)$ and $r$ is the smallest multiple of $k$ such that $v<r$ and $r-k\leq u\leq r$.
Proof.
First notice that this factorization always exists
$A=(A_{r})^{w}A_{v}=(A_{u}(s-1)^{r-u})^{w}A_{v}.$
If $w=1$ the $r=n$, $v=0$ and $A=A_{n}=A_{r}=A_{u}(s-1)^{r-u}$. Let ${\mathcal{B}}=\langle B,0\rangle$ be the pair obtained by undoing the transformation done by the $\theta$ operator, knowing that ${\mathcal{A}}$ and ${\mathcal{B}}$ are in $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$,
$\theta{\mathcal{B}}=\langle[B_{i-1}({b}_{i}-1)(s-1)^{j-i}]^{q}B_{n-qj},0\rangle={\mathcal{A}}$
where $i$ satisfies $b_{i+1}=\dots b_{n}=0$, and $i\leq j$ and $k|j$. Thus,
$[B_{i-1}(b_{i}-1)(s-1)^{j-i}]^{q}B_{n-qj}=(A_{u}(s-1)^{r-u})^{w}A_{v}.$
The word $B$ is determined by $r=j,u=i,w=q,v=n-qj$ and $A_{i-1}(a_{i}+1)=B_{i}$. Then,
$\displaystyle\langle B_{i-1}b_{i}0^{n-j},0\rangle=\langle A_{u-1}(a_{u}+1))0^{n-u},0\rangle,$
$\displaystyle\langle[B_{i-1}({b}_{i}-1)(s-1)^{j-i}]^{q}B_{n-qj},0\rangle=[A_{u-1}a_{u}(s-1)^{r-u}]^{w}A_{v}.$
Lemma 4.
If ${\mathcal{A}}\succ{\mathcal{B}}\succ{\mathcal{C}}$, with ${\mathcal{C}}={\mathcal{A}}\theta$, then ${\mathcal{B}}$ is not maximal .
Proof.
Let ${\mathcal{A}}=\langle A,0\rangle$, ${\mathcal{B}}=\langle B,0\rangle$, ${\mathcal{C}}={\mathcal{A}}\theta=\langle C,0\rangle$ and indices $i,j$ such that
$\displaystyle A$ $\displaystyle$== $\displaystyle{a}_{1}\cdots{a}_{i}0^{n-i},\text{with $\displaystyle a_{i}>0$},$
$\displaystyle B$ $\displaystyle$== $\displaystyle{b}_{1}\cdots{b}_{n},$
$\displaystyle C$ $\displaystyle$== $\displaystyle{c}_{1}\cdots{c}_{n}=[{a}_{1}\cdots{a}_{i-1}(a_{i}-1)(s-1)^{j-i}]^{q}{a}_{1}\cdots{a}_{n-qj}.$
Assume ${\mathcal{A}}\succ{\mathcal{B}}\succ{\mathcal{C}}$ and, by way of contradiction suppose ${\mathcal{B}}$ is a maximal pair. Since ${\mathcal{A}}\succ{\mathcal{B}}$,
$b_{1}=a_{1},\quad b_{2}=a_{2},\quad\ldots,\quad b_{i-1}=a_{i-1},$
and it should be $a_{i}>b_{i}$. Since ${\mathcal{B}}\succ{\mathcal{C}}$ then $b_{i}\geq c_{i}=a_{i}-1$; hence, $b_{i}=a_{i}-1$. Then, we have $B_{i}=C_{i}$. And for $\ell=1,\ldots j-i$ we have $b_{i+\ell}=s-1$. This is because ${\mathcal{B}}\succ{\mathcal{C}}$ and ${c}_{i+1}\cdots{c}_{i+\ell}=(s-1)^{j-i}$, then ${b}_{i+1}\cdots{b}_{i+\ell}=(s-1)^{j-i}$, given that $s-1$ is the lexicographically greatest symbol.
We now show that the other symbols in ${\mathcal{B}}$ and ${\mathcal{C}}$ also coincide. Since ${\mathcal{B}}$ is a maximal pair, $B\geq{b}_{j+1}\cdots{b}_{n}{b}_{1}\cdots{b}_{j}$ and we know that $a_{1}=b_{1}\geq b_{j+1}$. Since ${\mathcal{B}}\succ{\mathcal{C}}$ we have $b_{j+1}\geq c_{j+1}=a_{1}$, hence $b_{j+1}=a_{1}$. Repeating this argument we obtain for $m=1,..,q$ and $p=1,..,i-1$ we have:
$b_{mj+p}=a_{p}$, for $mj+p\leq n$,
$b_{mj+i}=a_{i}-1$ and
$b_{mj+i+\ell}=s-1$ for $\ell=1,..,j-i$.
This would imply ${\mathcal{B}}={\mathcal{C}}=\theta{\mathcal{A}}$. We conclude that there is no maximal pair ${\mathcal{B}}$ such that ${\mathcal{A}}\succ{\mathcal{B}}\succ{\mathcal{C}}$. ∎
Definition (list ${\mathcal{M}}$ of maximal pairs of $\Sigma^{n}\times\mathbb{Z}_{k}$).
Let $n$ and $k$ be positive integers such that $n|k$ or $k|n$. Define ${\mathcal{M}}$ by removing from $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$ the pairs that are not maximal. In case $n|k$ all the elements all maximal, none is removed. In case $k|n$ with $k<n$, for each pair ${\mathcal{A}}$ in $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$, except ${\mathcal{A}}=\langle 0^{n},0\rangle$, remove ${\mathcal{A}}\theta$, $A\theta^{2},...{\mathcal{A}}\theta^{h-1}$ where $h$ is the least such that $\theta^{h}({\mathcal{A}})$ is maximal . Let $M$ be the number of elements of the list ${\mathcal{M}}$.
Lemma 5.
The list ${\mathcal{M}}$ starts with the $\succ$-greatest pair $\langle(s-1)^{n},0\rangle$, ends with the $\succ$-smallest pair $\langle 0^{n},0\rangle$, and contains all maximal pairs in strictly decreasing $\succ$-order.
Proof of Lemma 5.
Case $k\mid n$.
  1. (1)
    By definition, the list ${\mathcal{M}}$ starts with $\langle(s-1)^{n},0\rangle$.
  2. (2)
    The maximal pairs are in $\succ$-decreasing order: This is because the list ${\mathcal{M}}$ is constructed by successive applications of the operator $\theta$ and by Lemma 2 the list is strictly $\succ$-decreasing.
  3. (3)
    No maximal pair is missing: Suppose ${\mathcal{A}}$ is in ${\mathcal{M}}$ and let $h$ be the least such that $\theta^{h}({\mathcal{A}})$ is maximal . To argue by contradiction, suppose there is a maximal pair ${\mathcal{B}}$ such that ${\mathcal{A}}\succ{\mathcal{B}}\succ\theta^{h}({\mathcal{A}})$. Then, there is $i$, $0\leq i<h$, such that
    ${\mathcal{A}}\succ\theta{\mathcal{A}}\succ\ldots\succ\theta^{i}{{\mathcal{A}}}\succ{\mathcal{B}}\succ\theta^{i+1}{{\mathcal{A}}}\succ\ldots\succ\theta^{h-1}{{\mathcal{A}}}\succ\theta^{h}{{\mathcal{A}}}.$
    Thus, ${\mathcal{B}}$ appears in between some ${\mathcal{D}}$ and $\theta{\mathcal{D}}$. By Lemma 4 this is impossible.
  4. (4)
    The list ${\mathcal{M}}$ ends with $\langle 0^{n},0\rangle$: There is no ${\mathcal{A}}$ such that $\langle 0^{n},0\rangle\succ{\mathcal{A}}$, and the list ${\mathcal{L}}$ is strictly $\succ$-decreasing. By Lemma 2, the operator $\theta$ applies to any pair except $\langle 0^{n},0\rangle$.
Case $n\mid k$, $n<k$: It is the same proof as in the previous case, but simpler because $\theta$ yields exactly all the maximal pairs in $\succ$-decreasing order. ∎
Example 1.
Let $s=2$, $n=6$ y $k=2$. The list ${\mathcal{M}}$ of maximal pairs is:
111111,0,111110,0,111101,0,111100,0,111010,0,111001,0,111000,0,110110,0,110101,0,110100,0,110010,0,110001,0,110000,0,101010,0,101001,0,101000,0,100101,0,100100,0,100001,0,100000,0,010101,0,010100,0,010000,0,000000,0111111011111001111010111100011101001110010111000011011001101010110100011001001100010110000010101001010010101000010010101001000100001010000000101010010100001000000000000$\langle$ 111111,0$\rangle$,$\langle$ 111110,0$\rangle$,$\langle$ 111101,0% $\rangle$,$\langle$ 111100,0$\rangle$,$\langle$ 111010,0$\rangle$,$\langle$ 111001,0% $\rangle$,$\langle$ 111000,0$\rangle$,$\langle$ 110110,0$\rangle$,\\ $\langle$ 110101,0$\rangle$,$\langle$ 110100,0$\rangle$,$\langle$ 110010,0$\rangle$,$\langle$ 110001,0$\rangle$,$\langle$ 110000,0$\rangle$,$\langle$ 101010,0$\rangle$,$\langle$ 101001,0$\rangle$,$\langle$ 101000,0$\rangle$,\\ $\langle$ 100101,0$\rangle$,$\langle$ 100100,0$\rangle$,$\langle$ 100001,0$\rangle$,$\langle$ 100000,0$\rangle$,$\langle$ 010101,0$\rangle$,$\langle$ 010100,0$\rangle$,$\langle$ 010000,0$\rangle$,$\langle$ 000000,0$\rangle$ 111111 , 0 ⟩ , ⟨ 111110 , 0 ⟩ , ⟨ 111101 , 0 ⟩ , ⟨ 111100 , 0 ⟩ , ⟨ 111010 , 0 ⟩ , ⟨ 111001 , 0 ⟩ , ⟨ 111000 , 0 ⟩ , ⟨ 110110 , 0 ⟩ , ⟨ 110101 , 0 ⟩ , ⟨ 110100 , 0 ⟩ , ⟨ 110010 , 0 ⟩ , ⟨ 110001 , 0 ⟩ , ⟨ 110000 , 0 ⟩ , ⟨ 101010 , 0 ⟩ , ⟨ 101001 , 0 ⟩ , ⟨ 101000 , 0 ⟩ , ⟨ 100101 , 0 ⟩ , ⟨ 100100 , 0 ⟩ , ⟨ 100001 , 0 ⟩ , ⟨ 100000 , 0 ⟩ , ⟨ 010101 , 0 ⟩ , ⟨ 010100 , 0 ⟩ , ⟨ 010000 , 0 ⟩ , ⟨ 000000 , 0 ⟩.
The next is the key lemma.
Lemma 6.
If ${\mathcal{A}}=\langle A,0\rangle$ is followed by ${\mathcal{B}}=\langle B,0\rangle$ in the list of maximal pairs ${\mathcal{M}}$ then $A$ is a prefix of $\overline{A}\overline{B}$.
Proof.
We can write ${\mathcal{A}}$ as
$\overline{{\mathcal{A}}}^{q}=\langle(A_{i}0^{p-i})^{q},0\rangle,$
where $q=n/p$ and $a_{i}>0$. If $q=1$ then $\overline{{\mathcal{A}}}={\mathcal{A}}=\langle A,0\rangle$ and $\overline{{\mathcal{A}}}\overline{{\mathcal{B}}}=\langle A,0\rangle\langle\overline{B},0\rangle=\langle A\overline{B},0\rangle$. Otherwise, $q>1$ and, since ${\mathcal{B}}=\theta{\mathcal{A}}^{h}$ for the smallest $h$ such that it is maximal , the shape of ${\mathcal{B}}$ is
${\mathcal{B}}=\langle\overline{A}^{q-1}A_{i-1}(a_{i}-1)(s-1)^{j-i}C,0\rangle$
for some word $C$ and for $j$ the smallest such that $i\leq j\leq p$ and $k\mid j$. Since ${\mathcal{B}}=\langle B,0\rangle$ and $B$ starts with $\overline{A}^{q-1}A_{i-1}(a_{i}-1)$ we have $\langle\overline{B},0\rangle=\langle B,0\rangle$, hence $\overline{{\mathcal{B}}}={\mathcal{B}}$. Then,
$\displaystyle\overline{{\mathcal{A}}}\overline{{\mathcal{B}}}=\langle\overline{A},0\rangle\langle\overline{B},0\rangle=\langle\overline{A},0\rangle\langle B,0\rangle=\langle\overline{A}B,0\rangle=\langle\overline{A}\overline{A}^{q-1}A_{i-1}(a_{i}-1)(s-1)^{j-i}C,0\rangle.$
In both cases $A$ is prefix of $\overline{A}\overline{B}$. ∎
3. Statement and proof of Theorem 1
Fix a finite alphabet $\Sigma$ with cardinality $s\geq 2$. Recall that a necklace is a circular word and a necklace is $(n,k)$-perfect if all words of length $n$ occur $k$ times but at positions with different congruence modulo $k$, for any convention of the starting position.
Theorem 1.
Let $n$ and $k$ be positive integers such that $k|n$ or $n|k$. The concatenation of the words in Lyndon pairs ordered lexicographically yields the lexicographically greatest $(n,k)$-perfect necklace.
Here is an example for $s=2$, $n=6$ y $k=2$. The lexicographically greatest $(n,k)$-perfect necklace is obtained by concatenating the following words (the symbol $\mid$ is just used here for ease of reading):
11111110111101111100111010111001111000110110110101110100110010110001110000101010011010001001011001001000011000000101010001000000conditional11delimited-∣∣111110111101delimited-∣∣111100111010delimited-∣∣111001111000delimited-∣∣110110110101delimited-∣∣110100110010delimited-∣∣110001110000delimited-∣∣10101001delimited-∣∣101000100101delimited-∣∣100100100001delimited-∣∣10000001delimited-∣∣01010001000000 11$\mid$ 111110$\mid$ 111101$\mid$ 111100$\mid$ 111010$\mid$ 111001$\mid$ 11% 1000$\mid$ 110110$\mid$ 110101$\mid$ 110100$\mid$ 110010$\mid$ 110001$\mid$ 110000$\mid$ 10% $\mid$ 101001$\mid$ 101000$\mid$ 100101$\mid$ 100100$\mid$ 100001$\mid$ 100000$\mid$ 01$\mid$ 0% 10100$\mid$ 010000$\mid$ 0011 ∣ 111110 ∣ 111101 ∣ 111100 ∣ 111010 ∣ 111001 ∣ 111000 ∣ 110110 ∣ 110101 ∣ 110100 ∣ 110010 ∣ 110001 ∣ 110000 ∣ 10 ∣ 101001 ∣ 101000 ∣ 100101 ∣ 100100 ∣ 100001 ∣ 100000 ∣ 01 ∣ $0$ $0$ 00
We nowgive the proof of Theorem 1. We start with a definition.
Definition (list ${\mathcal{L}}$ of Lyndon pairs).
Let $n$ and $k$ be positive integers such that $k|n$ or $n|k$. We define the list ${\mathcal{L}}$ as the list of Lyndon pairs of the elements in the list ${\mathcal{M}}$. Since ${\mathcal{M}}$ has $M$ elements, ${\mathcal{L}}$ has $M$ elements as well.
Assume $n|k$ or $k|n$. Let $X$ be the concatenation of the words in Lyndon pairs in the order given by ${\mathcal{L}}$.
3.1. Necklace $X$ has length $s^{n}k$
The length of $X$ is
$\sum_{\overline{{\mathcal{A}}}\text{ is Lyndon }}|\overline{{\mathcal{A}}}|.$
Each Lyndon pair is the reduction of a maximal pair. The length of a Lyndon pair ${\mathcal{A}}=\langle\overline{A},0\rangle$ is the length of the word $\overline{A}$, which is the number of different rotations of ${\mathcal{A}}$, see Observation 4. The necklace $X$ is defined by concatenating the words of all the reduced maximal pairs of $\Sigma^{n}\times\mathbb{Z}_{k}$ in $\succ$-decreasing order. Suppose $\overline{{\mathcal{A}}}=\langle\overline{A},0\rangle=\langle{a}_{1}\cdots{a}_{p},0\rangle$. Then ${\mathcal{A}}$ has $p$ different rotations : $\langle{a}_{1}\cdots{a}_{p},0\rangle$, $\langle{a}_{2}\cdots{a}_{n}a_{1},1\rangle$,…, $\langle a_{p}{a}_{1}\cdots{a}_{n-1},p-1\rangle$. So, each of the $p$ symbols of $\overline{A}$ is the start of a different rotation of ${\mathcal{A}}$.
By Lemma 6 we know that if ${\mathcal{A}}=\langle A,0\rangle$ is a maximal pair followed by the maximal pair ${\mathcal{B}}=\langle B,0\rangle$ then in $X$ we have $\overline{A}\overline{B}$. and $A$ is a prefix of $\overline{A}\overline{B}$. Since we concatenate in $X$ all the words of the reduced maximal pairs exactly once, we conclude that there is exactly one position for each of the different pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$. Thus, the length of $X$ is $s^{n}k$.
3.2. Necklace $X$ is $(n,k)$-perfect, with $k|n$
To prove that $X$ is a ($n$,$k$)-perfect necklace we need to show that each word of length $n$ occurs exactly $k$ times, at positions with different congruence modulo $k$. In this proof we number the positions of $X$ starting at 00 0; this is convenient for the presentation because the positions with congruence 00 0 are multiple of $k$.
We say that we find a pair $\langle A,u\rangle$ in $X=x_{0}x_{1}...x_{s^{n}k-1}$ when there is a position $p$ in $X$ such that $0\leq p<s^{n}k$ and $x_{p}..x_{p+|A|-1}=A$. To prove that $X$ is a $(n,k)$-perfect necklace we need to find all the rotations of all the maximal pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$ in the necklace $X$. Each maximal pair ${\mathcal{A}}=\langle A,0\rangle$ has eaxactly $|\overline{A}|$ many rotations.
Case $A=(s-1)^{n}$.
Since $k|n$, $\overline{{\mathcal{A}}}=\langle\overline{A},0\rangle=\langle(s-1)^{k},0\rangle$. Application of $\theta$ on ${\mathcal{A}}$ yields the maximal pair
$\theta{\mathcal{A}}={\mathcal{B}}=\langle B,0\rangle=\langle(s-1)^{n-1}(s-2),0\rangle.$
Since $\overline{{\mathcal{A}}}$ and $\overline{{\mathcal{B}}}$ are consecutive Lyndon pairs in ${\mathcal{L}}$, the construction of $X$ puts $\overline{A}$ followed by $\overline{B}$.
$\overline{{\mathcal{A}}}\overline{{\mathcal{B}}}=\langle(s-1)^{k},0\rangle\langle(s-1)^{n-1}(s-2),0\rangle=\langle(s-1)^{n+k-1}(s-2),0\rangle.$
This yields all rotations of ${\mathcal{A}}$: $\langle(s-1)^{n},0\rangle$, .., $\langle(s-1)^{n},k-1\rangle$.
Case $A=0^{n}$.
Since $k|n$, $\overline{{\mathcal{A}}}=\langle\overline{A},0\rangle=\langle 0^{k},0\rangle$. Consider the maximal pairs for $i=0,1,..k-1$.
${\mathcal{B}}=\langle B,0\rangle=\langle 0^{i}10^{k-1-i}0^{n-k},0\rangle.$
For each of these ${\mathcal{B}}$ the next maximal pair is
$\theta{\mathcal{B}}={\mathcal{C}}=\langle C,0\rangle=\langle(0^{i+1}(s-1)^{k-1-i})^{n/k},0\rangle.$
Since $\overline{{\mathcal{C}}}=\langle 0^{i+1}(s-1)^{k-1-i},0\rangle$, it is clear that ${\mathcal{C}}$ is maximal because all the rotations of ${\mathcal{C}}$ that have second component 00 0 are identical to ${\mathcal{C}}$. Since $\overline{{\mathcal{B}}}$ and $\overline{{\mathcal{C}}}$ are consecutive Lyndon pairs in ${\mathcal{L}}$, the construction of $X$ puts $\overline{B}$ followed by $\overline{C}$. Notice that for each $i=0,1,..,k-1$
¯$\overline{{\mathcal{B}}}\overline{{\mathcal{C}}}=\langle 0^{i}10^{k-1-i}0^{n-k}0^{i+1}(s-1)^{k-i-1},0\rangle=\langle 0^{i}10^{n}(s-1)^{k-i-1},0\rangle$
gives rise to the pair $\langle 0^{n},i+1\rangle$. Thus, we have all the rotations ${\mathcal{A}}=\langle 0^{n},0\rangle$, which are $\langle 0^{n},0\rangle$, .., $\langle 0^{n},k-1\rangle$.
Case $(s-1)^{n}>A>0^{n}$
Every maximal pair ${\mathcal{A}}=\langle A,0\rangle$ different from $\langle 0^{n},0\rangle$ has the form
${\mathcal{A}}=\langle(A_{p})^{q+1},0\rangle,$
where $A_{p}$ is $\overline{A}$ with $p$ the smallest integer such that $A=(A_{p})^{q+1}$ and $q+1=(n/p)$.
Subcase $q>0$. The pair ${\mathcal{A}}$, different from $\langle 0^{n},0\rangle$, always has a successor maximal pair in the list ${\mathcal{M}}$:
${\mathcal{B}}=\theta{\mathcal{A}}=\langle(A_{p})^{q}A_{i-1}(a_{i}-1)(s-1)^{j-i}{b}_{i+1}\cdots{b}_{p},0\rangle,$
where $i$ is the largest with $a_{i}>0$, $j$ is the smallest multiple of $k$ with $j\geq i$. Notice that $A_{p}\neq 0^{p}$. Since $\overline{{\mathcal{B}}}={\mathcal{B}}$,
$\overline{{\mathcal{A}}}\overline{{\mathcal{B}}}=\langle(A_{p})^{q+1}A_{i-1}(a_{i}-1)(s-1)^{j-i}{b}_{i+1}\cdots{b}_{p},0\rangle,$
and it also yields the first $i$ left rotations of ${\mathcal{A}}$, which are $\langle{a}_{r+1}\cdots{a}_{n}A_{r},r\rangle$, with $0\leq r<i$.
It remains to identify in our constructed necklace $X$ the $p-i$ right rotations of ${\mathcal{A}}=\langle(A_{p})^{q+1},0\rangle$. These are of the form
$\langle a_{p-i+1}...a_{p}A_{p}^{q}A_{p-i},i\rangle,$
To see this consider ${\mathcal{Q}}$ the predecessor of ${\mathcal{A}}$ by $\theta$,
$\theta{\mathcal{Q}}={\mathcal{A}}=\langle(A_{p})^{q+1},0\rangle$
where
${\mathcal{Q}}=\langle A_{p-1}(a_{p}+1)0^{pq},0\rangle.$
To see that ${\mathcal{Q}}$ is a maximal pair consider first ${\mathcal{R}}=\langle A_{p},0\rangle$ which, by Observation 2, is maximal. We now argue that $A_{p}$ has no proper suffix ${a}_{i}\cdots{a}_{p}$ which coincides with $A_{p-i+1}$. If there were such a prefix we could construct the rotation of ${\mathcal{R}}$ given by the pair ${\mathcal{S}}=\langle{a}_{i}\cdots{a}_{p}A_{i-1},0\rangle$ and one of the following would be true:
  • ${\mathcal{S}}={\mathcal{R}}$: but this is impossible by Observation 4.
  • ${\mathcal{R}}\succ{\mathcal{S}}$: Since ${\mathcal{R}}=\langle A_{p-i+1}{a}_{p-i+2}\cdots{a}_{p},0\rangle$, ${\mathcal{S}}=\langle a_{i}...a_{p}A_{i-1},0\rangle$ and we assumed ${a}_{i}\cdots{a}_{p}=A_{p-i+1}$, necessarily ${a}_{p-i+2}\cdots{a}_{p}>A_{i-1}$. But this contradicts that ${\mathcal{R}}$ is a maximal pair.
  • ${\mathcal{S}}\succ{\mathcal{R}}$: There is a rotation of ${\mathcal{R}}$ which is $\succ$-greater than ${\mathcal{R}}$, contradicting that ${\mathcal{R}}$ is a maximal pair.
We conclude that all suffixes ${a}_{i}\cdots{a}_{p}$ of $R$ are lexicographically smaller than $A_{p-i}$. Therefore, all suffixes of ${a}_{i}\cdots{a}_{p-1}(a_{p}+1)$ of $Q$ are lexicographically smaller than or equal to $A_{p-i}$. We already argue that ${\mathcal{Q}}$ is indeed maximal . For any rotation of ${\mathcal{Q}}$ of the form $\langle{a}_{i}\cdots{a}_{p-1}(a_{p}+1)0^{pq}{a}_{1}\cdots{a}_{i-1},0\rangle$, for $i>1$, we argued that ${a}_{i}\cdots{a}_{p-1}(a_{p}+1)$ is lexicographically smaller than or equal to $A_{p-i}$. Moreover, ${a}_{i}\cdots{a}_{p-1}(a_{p}+1)0^{pq}$ is lexicographically smaller than $Q=A_{p-1}(a_{p}+1)0^{pq}$. For any rotation of ${\mathcal{Q}}$ that starts with a suffix of $0^{pq}$ it can not be maximal , because for any $m$, $Q_{m}>0^{m}$, otherwise ${\mathcal{A}}$ would not be maximal .
We have the maximal successive pairs ${\mathcal{Q}}=\langle Q,0\rangle$, $\theta{\mathcal{Q}}={\mathcal{A}}=\langle A,0\rangle$ and ${\mathcal{B}}=\theta{\mathcal{A}}=\langle B,0\rangle$. From the arguments above, $\overline{Q}=Q$ and $\overline{B}=B$. Since ${\mathcal{A}}=\langle(A_{p})^{q+1},0\rangle$ and we assumed $q>1$, $A$ is not equal to $\overline{A}$. Then, $\overline{Q}\overline{A}\overline{B}=Q\overline{A}B$. Observe that the last $pq$ symbols of $Q$ followed by $\overline{A}$ followed by the first $pq$ symbols of $B$ give rise to
$\langle 0^{pq}(A_{p})^{q+1},0\rangle,$
where 00 0 is because $p$ is a multiple of $k$. We now identify in this pair the $p-i$ rotations of ${\mathcal{A}}$ to the right. Since ${\mathcal{A}}=\langle A,0\rangle$ where $A=(A_{i}0^{p-i})^{q+1}$, after $r$ rotations to the right of ${\mathcal{A}}$, for $r=0,1,..,p-i$, we obtain
$\langle 0^{r}(A_{p})^{q}A_{i}0^{p-i-r},-r\rangle.$
Subcase $q=0$. We need to see that for the maximal pairs ${\mathcal{A}}=\langle A,0\rangle$ where $A$ is reduced, that is $A=\overline{A}$, we can find all rotations of ${\mathcal{A}}$ in the constructed $X$.
If $A=0^{k-1}10^{n-k}$ then $\theta{\mathcal{A}}={\mathcal{B}}=\langle B,0\rangle=\langle 0^{n},0\rangle$, which is the last maximal pair in the list ${\mathcal{M}}$ and $\overline{B}=0^{k}$. Then,
$\overline{A}\overline{B}=0^{k-1}10^{n-k}0^{k}=0^{k-1}10^{n},$
and we can find $k$ left rotations of ${\mathcal{A}}$ which are of the form
$\langle 0^{k-1-r}10^{n-k+r},r\rangle\text{ for $\displaystyle r=0,1,..k-1$}.$
Now assume $A\neq 0^{k-1}10^{n-k}$. Let’s write
${\mathcal{A}}=\langle A_{i}0^{n-i},0\rangle,$
with $i$ such that $a_{i}>a_{i+1}=\ldots=a_{n}=0$, ${\mathcal{B}}=\langle B,0\rangle=\theta{\mathcal{A}}$ and ${\mathcal{C}}=\langle C,0\rangle=\theta{\mathcal{B}}=\theta^{2}{{\mathcal{A}}}$. Then ${\mathcal{B}}$ is of the form
${\mathcal{B}}=\langle A_{i-1}(a_{i}-1)(s-1)^{j-i}{b}_{j+1}\cdots{b}_{p},0\rangle,$
with $j$ is the least multiple of $k$ with $j\geq i$. Since $A=\overline{A}$, and by Lemma 6 $B$ is a prefix of $\overline{BC}$, then $AB$ is a prefix of $\overline{ABC}$, and we can find the first $i$ left rotations of ${\mathcal{A}}$ in it, which are of the form
$\langle{a}_{r+1}\cdots{a}_{n}A_{r},r\rangle,\text{ for $\displaystyle r=0,1,..,i-1$}.$
It remains to find $n-i$ right rotations of ${\mathcal{A}}$, which are of the form
$\langle 0^{r}A_{i}0^{n-i-r},-r\rangle$
for $r=1,\ldots,n-i$. Equivalently we can write it as
$\langle 0^{n-i-h}A_{i}0^{h},(n-i-h)\rangle$
for $h=0,1,\ldots,n-i-1$.
If $h=0$ and $A_{i}=(s-1)^{i}$ then the pair is $\langle 0^{n-i}(s-1)^{i},n-i\rangle$. We find it in the pair $\langle 0^{n}(s-1)^{n},0\rangle$, which results from the concatenation of the words in the last two Lyndon pairs in ${\mathcal{L}}$ and the first two, that is ${\mathcal{L}}_{M-1}$ and ${\mathcal{L}}_{M}$, followed by ${\mathcal{L}}_{1}$ and ${\mathcal{L}}_{2}$.
If $h\neq 0$ or $A_{i}\neq(s-1)^{i}$ then we find the right rotations of ${\mathcal{A}}$ in the concatenation of the words of three Lyndon pairs, that we call $\overline{{\mathcal{Q}}_{t}}\overline{{\mathcal{P}}}\overline{{\mathcal{R}}}$ and we define below. Recall that ${\mathcal{A}}=\langle A,0\rangle=\langle A_{i}0^{n-i},0\rangle$, with $i$ such that $a_{i}>a_{i+1}=\ldots=a_{n}=0$, is a maximal pair. We claim there is a unique pair ${\mathcal{Q}}$ of the form
${\mathcal{Q}}=\langle A_{i}0^{h}S,0\rangle,$
where $|S|=n-(i+h)$ is not zero. Notice that ${\mathcal{Q}}$ may not be maximal.
Since ${\mathcal{Q}}$ is a pair $\succ$-greater than ${\mathcal{A}}$, we know ${\mathcal{Q}}$ is a predecessor of ${\mathcal{A}}$ by $\theta$. The pair ${\mathcal{Q}}_{t}$ is the closest predecessor of ${\mathcal{Q}}$ by $\theta$ that is a maximal pair. For this, consider the succesive application of the operator $\theta$, which allows us to traverse the list $(\theta^{l}\langle(s-1)^{n},0\rangle)_{1\leq l\leq T}$. Our interest is to traverse it backwards. Here is a diagram:
${\mathcal{Q}}_{t}$${\mathcal{Q}}\!=\!\langle A_{i}0^{h}S,0\rangle$${\mathcal{A}}$apply many times $\theta$apply many times $\theta$
For $l=1,2,3,...$ let ${\mathcal{Q}}_{l}=\langle Q_{l},0\rangle$ be the pair such that $\theta^{l}{{\mathcal{Q}}_{l}}={\mathcal{Q}}$. For $l=1$,
${\mathcal{Q}}_{1}=\langle Q_{1},0\rangle=\langle A_{u_{1}-1}(a_{u_{1}}+1)0^{n-u_{1}},0\rangle,$
$\theta{\mathcal{Q}}_{1}=\langle(A_{u_{1}}(s-1)^{r_{1}-u_{1}})^{w_{1}}A_{v_{1}}S,0\rangle=\langle(A_{r_{1}})^{w_{1}}A_{v_{1}}S,0\rangle=\langle A_{i}0^{h}S,0\rangle={\mathcal{Q}},$
where $r_{1}$ is the least multiple of $k$ such that $v_{1}<r_{1}$, $a_{u_{1}}<(s-1)$, $A_{r_{1}}=A_{u_{1}}(s-1)^{r_{1}-u_{1}}$ and $r_{1}-k\leq u_{1}\leq r_{1}$. If ${\mathcal{Q}}_{1}$ is a maximal pair then fix $t=1$ and we have finished the search. Otherwise we consider the predecessor of ${\mathcal{Q}}_{1}$ by $\theta$,
${\mathcal{Q}}_{2}=\langle Q_{2},0\rangle=\langle A_{u_{2}-1}(a_{u_{2}}+1)0^{n-u_{2}},0\rangle$
such that
$\theta{\mathcal{Q}}_{2}=\langle(A_{u_{2}}(s-1)^{r_{2}-u_{2}})^{w_{2}}A_{v_{2}},0\rangle=\langle(A_{r_{2}})^{w_{2}}A_{v_{2}},0\rangle={\mathcal{Q}}_{1},$
where $r_{2}$ is the least multiple of $k$, $v_{2}<r_{2}$, and $a_{u_{2}}<(s-1)$ such that $A_{r_{2}}=A_{u_{2}}(s-1)^{r_{2}-u_{2}}$. We know that $a_{u_{2}}<(s-1)$ exists because ${\mathcal{Q}}_{1}\neq\langle(s-1)^{n},0\rangle$ (otherwise ${\mathcal{Q}}_{1}$ would be a maximal pair), and then there is $a_{i}<s-1$ for some $1\leq i\leq r_{2}$. Notice that $u_{2}<u_{1}$. If ${\mathcal{Q}}_{2}$ is a maximal pair then we fix $t=2$ and we have finished the search. Otherwise, we repeat this procedure. In this way the predecessor of ${\mathcal{Q}}_{l-1}$ by $\theta$ is
${\mathcal{Q}}_{l}=\langle Q_{l},0\rangle=\langle A_{u_{l}-1}(a_{u_{l}}+1)0^{n-u_{l}},0\rangle,$
such that
$\theta{\mathcal{Q}}_{l}=\langle(A_{u_{l}}(s-1)^{r_{l}-u_{l}})^{w_{l}}A_{v_{l}},0\rangle=\langle(A_{r_{l}})^{w_{l}}A_{v_{l}},0\rangle={\mathcal{Q}}_{l-1},$
where $r_{l}$ is multiple of $k$, $v_{l}<r_{l}$, $u_{l}<u_{l-1}$ and $a_{u_{l}}<(s-1)$ such that
$A_{r_{l}}=A_{u_{l}}(s-1)^{r_{l}-u_{l}}.$
Eventually we find $t$ such that ${\mathcal{Q}}_{t}$ is a maximal pair.
Consider now the three consecutive maximal pairs in the list ${\mathcal{M}}$,
${\mathcal{Q}}_{t}=\langle Q_{t},0\rangle,{\mathcal{P}}=\langle P,0\rangle,\text{ and }{\mathcal{R}}=\langle R,0\rangle.$
Let $\overline{{\mathcal{Q}}_{t}},\overline{{\mathcal{P}}}$ and $\overline{{\mathcal{R}}}$ be the corresponding Lyndon pairs. ${\mathcal{P}}$ always exists, because it’s either ${\mathcal{A}}$ or a maximal pair before it, and ${\mathcal{R}}$ is $\langle 0^{n},0\rangle$ in the worst case. Notice that $\overline{Q_{t}}$ ends with $0^{n-u_{t}}$ and, by Lemma 6, $\overline{P}\overline{R}$ starts with $P$. Therefore, $\overline{{\mathcal{Q}}_{t}}\overline{{\mathcal{P}}}\overline{{\mathcal{R}}}$ contains $0^{n-u_{t}}P=0^{n-u_{t}}A_{i}0^{h}C$, for some $C$. Since ${\mathcal{P}}$ is a maximal pair $\succ$-greater than or equal to ${\mathcal{A}}$, we can assert that its prefix is $A_{i}0^{h}$. Finally, we have
$u_{t}\leq u_{1}\leq i+h,$
because $u_{t}<u_{t-1}<\ldots<u_{1}$ and $u_{1}\leq i+h$ because $u_{1}$ was the position of a symbol in $A_{i}0^{h}$, which has length $i+h$. Then,
$\displaystyle u_{t}\leq u_{1}\leq i+h$ $\displaystyle\iff-u_{t}\geq-u_{1}\geq-i-h$
$\displaystyle\iff n-u_{t}\geq n-u_{1}\geq n-i-h$
$\displaystyle\iff n-u_{t}\geq n-i-h.$
We conclude that $\overline{{\mathcal{Q}}_{t}}\overline{{\mathcal{P}}}\overline{{\mathcal{R}}}$ contains
$\langle 0^{n-i-h}A_{i}0^{h},(|\overline{Q_{h}}|-(n-i-h))\rangle.$
This can be rewritten as the pair
$\langle 0^{n-i-h}A_{i}0^{h},-(n-i-h)\rangle,$
because $|\overline{Q_{t}}|$ is a multiple of $k$, hence ${\mathcal{Q}}_{t}$ has second component 00 0.
3.3. Necklace $X$ is $(n,k)$-perfect, with $n|k$, $n<k$
This proof is similar to the the proof for the case $k|n$, but it uses the corresponding definitions of $\theta$, the list ${\mathcal{M}}$ and the notion of expansion for the list ${\mathcal{L}}$. The proof becomes simpler because it requires fewer cases and each case is easier. To find the rotations of $\langle 0^{n},0\rangle$ is similar to the case $k|n$. To find the rotations to the left for pairs $\langle A,0\rangle$ such that $(s-1)^{n}>A>0^{n}$ is also similar to the case $k|n$, because the number of positions $i$ to be considered in the two definitions of $\theta$ coincide. However, to find rotations to the right now is much simpler than in the case $k|n$ because the operator $\theta$ is now a bijection. Given that all the pairs obtained by application of $\theta$ are maximal pairs, for each pair we just need to find the predecessor by $\theta$. In contrast, when $k|n$, the operator $\theta$ is not bijective and we had to justify how we find the maximal pair ${\mathcal{Q}}_{t}$.
3.4. Necklace $X$ is the lexicographically greatest $(n,k)$-perfect necklace
Given an $(n,k)$-perfect necklace fix the starting position 00 0 in the first symbol of the occurrence of $(s-1)^{n}$ that yields the lexicographically maximal rotation $a_{0}\ldots a_{s^{n}k-1}$.. For each position $i$ such that $1\displaystyle 0\leq i\leq ks^{n}-$ 1 define the pair $\langle a_{i}...a_{i+n-1},i\mod k\rangle$. This of course includes the pairs starting at positions $s^{n}k-n+1$, …, $s^{n}k-1$ because we are dealing with a necklace. Notice that the symbol $a_{i+n-1}$ is not just in the pair starting at position $i$, but also also in the pairs starting at positions $i+1,...,i+n-1$. The definition of $(n,k)$-perfect necklace ensures that this set of pairs is exactly $\Sigma^{n}\times\mathbb{Z}_{k}$. In the lexicographically greatest $(n,k)$-perfect necklace, as $i$ increases from 00 0 to $s^{n}k-1$, at each position $i$ we find the $\succ$-greatest remaining pair $\langle a_{i}...a_{i+n-1},i\mod k\rangle$ given that the symbols $a_{i},..,a_{i+n-2}$ have already been determined. Thus, the choices for the maximal pair starting at position $i$ are just the choices for the symbol $a_{i+n-1}$. Let’s see how our construction achieves this.
Recall that $M$ is the number of maximal pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$ with second component 00 0. This is exactly the length of the list ${\mathcal{M}}$ and the length of the list ${\mathcal{L}}$. Our construction concatenates all the words in Lyndon pairs of $\Sigma^{n}\times\mathbb{Z}_{k}$ in $\succ$-decreasing order, given in ${\mathcal{L}}=(\langle w_{i},0\rangle)_{1\leq i\leq M}$. So, ${\mathcal{L}}$ is the list of all the reduced maximal pairs of $\Sigma^{n}\times\mathbb{Z}_{k}$ with second component 00 0 in $\succ$-decreasing order. Let $\ell_{0}=0$ and $(\ell)_{1\leq i\leq M}$ be sum of the lengths of the first $i$ Lyndon pairs in the list ${\mathcal{L}}$,
$\ell_{i}=\sum_{j=1}^{i}|{\mathcal{L}}_{j}|.$
As we already showed, $\ell_{M}=s^{n}k$. Our construction yields the word $X=a_{0},a_{1},\ldots a_{s^{n}k-1}$ where ${\mathcal{L}}_{1}=\langle a_{0},\ldots a_{\ell_{1}-1},0\rangle$, ${\mathcal{L}}_{2}=\langle a_{\ell_{1}},\ldots a_{\ell_{2}-1},0\rangle$, …, ${\mathcal{L}}_{M}=\langle a_{\ell_{M-1}},\ldots,\ldots a_{\ell_{M}-1},0\rangle$. That is, for each $i$ such that $1\displaystyle 0\leq i\leq s^{n}k-$ 1, ${\mathcal{L}}_{i}=\langle w_{i},0\rangle,\text{ where }w_{i}=a_{\ell_{i-1}},\ldots a_{\ell_{i}-1}$, $w_{i}$ is lexicographically greater than all of its rotations, and $w_{i}$ is lexicographically greater than all the rotations of all the $w_{j}$, for $j>i$. Our construction greedily picks, one after the other, the $\succ$-greatest remaining reduced maximal pair with second component 00 0. By Lemma 6 we know that the concatenation of the reduced maximal pairs in decreasing $\succ$-order gives rise to the same pairs without reduction.
Implicitly our construction defines a function $f:\Sigma^{n}\to\Sigma$, such that
$a_{i+n}=f(a_{i},a_{i+1},...,a_{i+n-1}).$
Since the $\succ$-order among the pairs with the same second component is just the lexicographical order, $a_{0},a_{2},\ldots a_{s^{n}k-1}$ is the lexicographically greatest among all $(n,k)$-perfect necklaces because in position $i+n$ the symbol $a_{i+n}$ is the lexicographically greatest among all the symbols that could have been allocated to define a perfect necklace. Consequently, if $b_{0},b_{1},\ldots,b_{s^{n}k-1}$ is another $(n,k)$-perfect necklace then, necessarily, there is a position $p$ such that $b_{0}=a_{0},b_{1}=a_{1},\ldots,b_{p-1}=a_{p-1}$ but $a_{p}>b_{p}$. This completes the proof of Theorem 1.
Acknowledgements
This work is supported by the projects CONICET PIP 11220210100220CO and UBACyT 20020220100065BA.