### v2

# Lyndon pairs and the lexicographically greatest perfect necklace

Abstract.
Fix a finite alphabet. A necklace is a circular word. For positive integers $n$ and $k$, a necklace is $(n,k)$-perfect if all words of length $n$ occur $k$ times but at positions with different congruence modulo $k$, for any convention of the starting position. We define the notion of a Lyndon pair and we use it to construct the lexicographically greatest $(n,k)$-perfect necklace, for any $n$ and $k$ such that $n$ divides $k$ or $k$ divides $n$. Our construction generalizes Fredricksen and Maiorana’s construction of the lexicographically greatest de Bruijn sequence of order $n$, based on the concatenation of the Lyndon words whose length divide $n$.
Key words and phrases:
de Bruijn sequences, Lyndon words
2020 Mathematics Subject Classification:
Primary 68R15, 05A05; Secondary 11K16
1. Introduction
Let $\Sigma$ be a finite alphabet with at least two symbols. A word on $\Sigma$ is a finite sequence of symbols, and a necklace is the equivalence class of a word under rotations. Given two positive integers, $n$ and $k$, a necklace is $(n,k)$-perfect if all words of length $n$ occur $k$ times but at positions with different congruence modulo $k$, for any convention of the starting position. The well known circular de Bruijn sequences of order $n$, see [3, 7, 8], are exactly the $(n,k)$-perfect necklaces for $k=1$. For example, 1110010011100100 1110010011100100 is a $(2,2)$-perfect for $\Sigma=\{0,1\}$. The $(n,k)$-perfect necklaces correspond to Hamiltonian cycles in the tensor product of the de Bruijn graph with a simple cycle of length $k$.
A thorough presentation of perfect necklaces appears in [1]. With the purpose of constructing normal numbers with very fast convergence to normality M. Levin in [9] gives two constructions of perfect necklaces. One based on arithmetic progressions with difference coprime with the alphabet size which yields $(n,n)$-perfect necklaces. The other based on Pascal triangle matrix which yields nested $(n,n)$-perfect necklaces when $n$ is a power of 22 22. In [2] there is a method of constructing all nested $(n,n)$-perfect necklaces for the alphabet $\{0,1\}$.
Assume the lexicographic order on words. A Lyndon word is a nonempty aperiodic word that is lexicographically greater than all of its rotations. For example, the Lyndon words over alphabet $\{0,1\}$ sorted by length and then in decreasing lexicographical order within each length are
1,0,10,110,100,1110,1100,1000,11110,11100,11010,11000,10100,10000,1010110100111011001000111101110011010110001010010000 1,$0$$\ldots$1 , 0 , 10 , 110 , 100 , 1110 , 1100 , 1000 , 11110 , 11100 , 11010 , 11000 , 10100 , 10000 , …
Lyndon words were introduced by Lyndon in the 1950s [10, 11]. They provide a nice factorization of the free monoid $\Sigma^{*}$: each word $w$ in $\Sigma^{*}$ has a unique decomposition as a product $w=u_{1}\ldots u_{n}$ of a non-increasing sequence of Lyndon words $u_{1},\ldots u_{n}$ in the lexicographic order. The problem to compute the prime factorization of a given word has a solution in time linear to the length of the given word [4], see also [6].
Fredricksen and Maiorana [5] construct a de Bruijn sequence of order $n$ by concatenating all the Lyndon words whose length divides $n$. They first identify each necklace with the word that represents the lexicographically maximal rotation. Order the necklaces according to lexicographical order of these words. Fredericksen and Maiorana de Bruijn sequence of order $n$ is the concatenation, according to this order of the necklaces, of the respective Lyndon words having length divisible by $n$. For example, the binary words of length 44 44 yield 66 66 necklaces, and the words representing the lexicographically maximal rotations, in decreasing lexicographical order are:
$1$, 1110, 1100, 1010, 1000, $0$.111111101100101010000000 $1$,\ \ 1110,\ \ 1100,\ \ 1010,\ \ 1000,\ \ $0$.1111 , 1110 , 1100 , 1010 , 1000 , $0$
The corresponding Lyndon words, in the above order are:
1, 1110,1100, 10, 1000, 0.111101100101000 0 1,\ \ 1110,1100,\ \ 10,\ \ 1000,\ \ 0.1 , 1110 , 1100 , 10 , 1000 , 0 .
The length of each of them divides $n=4$, because it is 11 11 or 22 22 or 44 44. Then, none of them is discarded, hence, Fredericksen and Maiorana’s de Bruijn sequence of order $n$ sequence is:
1 1110 1100 10 1000 0.1111011001010000 1\ 1110\ 1100\ 10\ 1000\ 0.1 1110 1100 10 1000 0 .
This construction, together with the efficient generation of Lyndon words, provides a method for constructing the lexicographically greatest de Bruijn necklace of each order $n$ in linear time and logarithmic space.
In this note we present the notion of Lyndon pairs and we use it to generalize Fredricksen and Maiorana’s algorithm to construct the lexicographically greatest $(n,k)$-perfect necklaces, for any $n$ and $k$ such that $n$ divides $k$ or $k$ divides $n$. Our presentation also elaborates the rather briefly presented ideas and proofs in [5].
2. Lyndon pairs in lexicographical order
2.1. Lyndon pairs
We assume a finite alphabet $\Sigma$ with cardinality $s$, with $s\geq 2$. Without loss of generality we assume $\Sigma=\{0,\ldots,s-1\}$. We use lowercase letters $a,b,c$ possibly with subindices for alphabet symbols. Words are finite sequences of symbols that we write $a_{1}a_{2}\ldots a_{n}$ or with a capital letter $A,B,C$. We write $a^{\ell}$ to denote the word of length $\ell$ made just of $a^{\prime}$s. and we write $A^{\ell}$ to denote the word made of $\ell$ copies of $A$. The concatenation of two words $A$ and $B$ is written $AB$. The length of a word $A$ is denoted with $|A|$. The positions of a word $A$ are numbered from 11 11 to $|A|$. We use $>$ to denote the decreasing lexicographic order on words and we write $A\geq B$ when $A>B$ or $A=B$.
We use lowercase letters $h,\ldots,z$ to denote non-negative integers. We write $k|n$ to say that $k$ divides $n$. We write $\mathbb{Z}_{k}$ for the set $\mathbb{Z}/k\mathbb{Z}$ of residues modulo $k$. We also use $<$ and $>$ for the natural orders on $\mathbb{Z}_{k}$ and $\mathbb{N}$ and, as usual, $u\leq v$ when $u<v$ or $u=v$; and $v\geq u$ when $v>u$ or $v=u$. When $u<v$ we may write $v>u$.
We work with pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$ when $k|n$ or $n|k$. This condition on $n$ and $k$ is assumed all along the sequel. We refer to pairs $\langle A,u\rangle$ with calligraphic letter ${\mathcal{A}},{\mathcal{B}},{\mathcal{C}},\ldots$. If ${\mathcal{A}}=\langle A,u\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ we write ${\mathcal{A}}^{p}$ to denote $\langle A^{p},u\rangle$ in $\Sigma^{pn}\times\mathbb{Z}_{k}$ We consider the following order $\succ$ over $\Sigma^{n}\times\mathbb{Z}_{k}$.
Definition (order $\succ$ on $\Sigma^{n}\times\mathbb{Z}_{k}$).
$\langle A,u\rangle\succ\langle B,v\rangle\text{ exactly when either }(A>B\text{ and }u=v)\text{ or }(k-1-u)>(k-1-v).$
The smallest the second component in $\mathbb{Z}_{k}$, the $\succ$-greater the pair. Among pairs with the same second component in $\mathbb{Z}_{k}$, the order $\succ$ is defined with the decreasing lexicographic order on $\Sigma^{n}$. Thus, $\langle(s-1)^{n},0\rangle$ is the $\succ$-greatest in $\Sigma^{n}\times\mathbb{Z}_{k}$ and $\langle 0^{n},k-1\rangle$ is the $\succ$-least. As usual, we write $\langle A,u\rangle\succeq\langle B,v\rangle$ exactly when $\langle A,u\rangle\succ\langle B,v\rangle$ or $\langle A,u\rangle=\langle B,v\rangle$.
Definition (rotation of a pair).
Given a pair $\langle a_{1}\ldots a_{n},u\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ its right rotation is the pair $\langle a_{n}a_{1}\ldots a_{n-1},u-1\rangle$ and its left rotation is the pair $\langle a_{2}\ldots a_{n}a_{1},u+1\rangle$.
For $s=3$, $n=k=5$ and $\langle 13212,4\rangle$, its right rotation is $\langle 21321,3\rangle$, and its left rotation is $\langle 32121,0\rangle$. The rotation function induces a relation between pairs: two pairs are related if successive rotations initially applied to the first yield the second. This relation is clearly reflexive and transitive. For pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$, when $k|n$ or $n|k$, the rotation has an inverse, given by successive rotations. So the relation is also symmetric, hence, an equivalence relation.
Definition (maximal pair).
A necklace in $\Sigma^{n}\times\mathbb{Z}_{k}$ is a set of pairs that are equivalent under rotations. For each $u=0,\ldots k-1$, the following set is a necklace
$\{\langle a_{i+1}\ldots a_{n}a_{1}\ldots a_{i},u+i\rangle:0\leq i<\max(n,k)\}.$
In each necklace we are interested in the pair that is maximal in the order $\succ$. If a pair ${\mathcal{A}}$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ is $\succeq$-maximal among its rotations we call it maximal.
For example, for $n=4$ and $k=2$ the pair $\langle 1110,0\rangle$ is maximal because it is $\succ$-greater than its rotations $\langle 1101,1\rangle$, $\langle 1011,0\rangle$ and $\langle 0111,1\rangle$. The pair $\langle 1010,0\rangle$ is maximal because it is $\succ$-greater than its rotation $\langle 0101,1\rangle$.
Observation 1.
When $n|k$ the maximal pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$ are the pairs $\langle A,0\rangle$ for $A\in\Sigma^{n}$.
We can concatenate pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$ having the same second component: the concatenation of $\langle A,u\rangle$ and $\langle B,u\rangle$ is $\langle AB,u\rangle$. Given ${\mathcal{A}}=\langle A,u\rangle$ we write ${\mathcal{A}}^{p}$ to denote the pair $\langle A^{p},u\rangle$.
Observation 2.
Assume $k|n$ or $n|k$. Then, ${\mathcal{A}}\in\Sigma^{n}\times\mathbb{Z}_{k}$ is maximal exactly when, for any $p\geq 1$, ${\mathcal{A}}^{p}$ is maximal in $\Sigma^{pn}\times\mathbb{Z}_{k}$.
Proof.
($\implies$). If $n|k$ then all pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$ with second component 00 0 are maximal . We prove it for $k|n$, $k<n$. Assume ${\mathcal{A}}=\langle A,u\rangle$ such that ${\mathcal{A}}^{p}=\langle A^{p},u\rangle$ is maximal , but ${\mathcal{A}}$ is not maximal. Then ${\mathcal{A}}$ has a rotation with $j\geq 1$ such that ${\mathcal{R}}=\langle({a}_{j+1}\cdots{a}_{n}{a}_{1}\cdots{a}_{j}),u+j\rangle$ and ${\mathcal{R}}\succ{\mathcal{A}}$. Then either,
${a}_{j+1}\cdots{a}_{n}{a}_{1}\cdots{a}_{j}>{a}_{1}\cdots{a}_{n},\text{ or }(k-1-(u+j))>(k-1-u).$
But this implies that either
$({a}_{j+1}\cdots{a}_{n}{a}_{1}\cdots{a}_{j})^{p}>({a}_{1}\cdots{a}_{n})^{p},\text{ or }(k-1-(u+j))>(k-1-u).$
This, in turn, implies ${\mathcal{R}}^{p}\succ{\mathcal{A}}^{p}$, which contradicts that ${\mathcal{A}}^{p}$ is maximal.
[…]
${\mathcal{R}}=\langle R,0\rangle=\langle{a}_{1+j}\cdots{a}_{n}{a}_{1}\cdots{a}_{j}A^{p-1},0\rangle.$
Given that the second component of both ${\mathcal{R}}$ and ${\mathcal{A}}^{p}$ is 00 0, necessarily $R>A^{p}$. But this contradicts that ${\mathcal{A}}=\langle A,0\rangle$ was a maximal rotation. ∎
Observation 3.
If $k|n$ and ${\mathcal{A}}\in\Sigma^{n}\times\mathbb{Z}_{k}$ is a maximal pair then none of its rotations are $\succ$-greater than ${\mathcal{A}}$, but there may be a rotation that is equal to ${\mathcal{A}}$.
For a word $A=a_{1}\ldots a_{n}$ we write $A_{i}$ to denote its prefix of length $i$, that is, $a_{1}\ldots a_{i}$.
Lemma 1.
Let $n$ and $k$ be positive integers such that $k|n$ or $n|k$. Let ${\mathcal{A}}=\langle A,0\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ be maximal and different from $\langle 0^{n},0\rangle$. Suppose $A=a_{1}\ldots a_{n}$, let $i$ be such that $a_{i}>0$ and let
${\mathcal{B}}=\langle A_{i-1}(a_{i}-1)(s-1)^{j-i},0\rangle,$
where if $n|k$ then $j=n$; and if $k|n$ then $j$ is the smallest such that $k|j$ and $i\leq j<n$. Then, ${\mathcal{B}}\in\Sigma^{j}\times\mathbb{Z}_{k}$ is maximal.
Proof.
Let ${\mathcal{A}}$ be a maximal pair . By way of contradiction, assume ${\mathcal{B}}$=$\langle A_{i-1}(a_{i}-1)(s-1)^{j-i},0\rangle$ is not a maximal pair. Then there is some $\ell$ multiple of $k$ such that :
$\langle{a}_{\ell+1}\cdots{a}_{i-1}(a_{i}-1)(s-1)^{j-i}A_{\ell},0+\ell\rangle\succ\langle A_{i-1}(a_{i}-1)(s-1)^{j-i},0\rangle$
Necessarily
${a}_{\ell+1}\cdots{a}_{i-1}\geq{a}_{1}\cdots{a}_{(i-1)-(\ell+1)+1}.$
Since ${\mathcal{A}}$ is a maximal pair,
${a}_{1}\cdots{a}_{(i-1)-(\ell+1)+1}\geq{a}_{\ell+1}\cdots{a}_{i-1}.$
Therefore, $a_{\ell+1}=a_{1}$, $a_{\ell+2}=a_{2}$, $\ldots$, $a_{i-1}=a_{(i-1)-(\ell+1)+1}$. This implies, $a_{i}-1\geq a_{(i-1)-(\ell+1)+2}$. Consequently,
${a}_{\ell+1}\cdots{a}_{i-1}a_{i}>{a}_{1}\cdots{a}_{(i-1)-(\ell+1)+2},$
but this contradicts that ${\mathcal{A}}$ is a maximal pair. ∎
For example, let $s=7$, $n=6$, $k=3$ and ${\mathcal{A}}=\langle 456123,0\rangle$. Since ${\mathcal{A}}$ is maximal and the symbol in position $i=3$ is 66 66, letting $j=3$ (a multiple of $k$ grater than or equal to $i$) we obtain in $\Sigma^{j}\times\mathbb{Z}_{k}$ the maximal pair ${\mathcal{B}}=\langle 455,0\rangle$. No filling with $s-1=6$ was required. Since the symbol in position $i=1$ is 44 44, letting $j=3$ we obtain the maximal pair ${\mathcal{B}}=\langle 366,0\rangle$ which is obtained by concatenating $s-1=6$ $(j-i)=2$ times.
Definition (reduction).
Let $n$ and $k$ be positive integers such that $k|n$. The reduction of a word $A={a}_{1}\cdots{a}_{n}$ is the word $\overline{A}={a}_{1}\cdots{a}_{p}$ where $({a}_{1}\cdots{a}_{p})^{n/p}={a}_{1}\cdots{a}_{n}$ and $p$ is the smallest such that $k\mid p$, $p\mid n$. The reduction of a pair ${\mathcal{A}}=\langle A,u\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$, is the pair $\overline{{\mathcal{A}}}=\langle\overline{A},u\rangle$ in $\Sigma^{p}\times\mathbb{Z}_{k}$.
When $k|n$, the reduction always exists because one can take $p=n$. Notice that all the rotations of a reduced pair are pairwise different. For example, for $s=8$, $n=8$ and $k=2$, $\overline{\langle 10101010,0\rangle}=\langle 10,0\rangle$; $\overline{\langle 01230123,0\rangle}=\langle 0123,0\rangle$; $\overline{\langle 01234567,0\rangle}=\langle 01234567,0\rangle$.
Definition (expansion).
Let $n$ and $k$ be positive integers such that $n|k$. The expansion of a word $A={a}_{1}\cdots{a}_{n}$ is the word $\widetilde{A}=({a}_{1}\cdots{a}_{n})^{k/n}$. The expansion of a pair ${\mathcal{A}}=\langle A,u\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ is the pair $\widetilde{{\mathcal{A}}}=\langle\widetilde{A},u\rangle$ in $\Sigma^{k}\times\mathbb{Z}_{k}$.
When $n|k$ the expansion always exists. For example for $s=3$, $n=2$ and $k=8$, $\widetilde{\langle 12,0\rangle}=\langle 12121212,0\rangle$. Notice that when $k=n$ then $\widetilde{A}=A$.
Observation 4.
If $n|k$ then for every pair ${\mathcal{A}}=\langle A,u\rangle\in\Sigma^{n}\times\mathbb{Z}_{k}$, $A$ is a prefix of $\widetilde{A}$.
Definition (Lyndon pair).
Let $n$ and $k$ be positive integers. When $k|n$, the Lyndon pairs are the reductions of the maximal pairs in $\Sigma^{n}\times{\mathbb{Z}_{k}}$. When $n|k$ with $n<k$, the Lyndon pairs are the expansions of the maximal pairs $\langle A,0\rangle\in\Sigma^{n}\times\mathbb{Z}_{k}$.
Thus, when $k|n$, the Lyndon pairs are elements in $\bigcup_{p:k|p|n}\Sigma^{p}\times\mathbb{Z}_{k}$. But when $n|k$ the Lyndon pairs are elements in $\Sigma^{k}\times\mathbb{Z}_{k}$.
Observation 5.
Every Lyndon pair is strictly $\succ$-greater than each of its rotations.
2.2. The operator $\theta$
We define the operator $\theta$ that given a pair in $\Sigma^{n}\times\mathbb{Z}_{k}$ different from $\langle 0^{n},0\rangle$ but with second component 00 0, it defines another pair in $\Sigma^{n}\times\mathbb{Z}_{k}$ with second component 00 0.
Definition (operator $\theta$).
Let $n$ and $k$ be positive integers such that $k|n$ or $n|k$. For ${\mathcal{A}}=\langle A,0\rangle=\langle{a}_{1}\cdots{a}_{n},0\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ such that $a_{i}>a_{i+1}=\ldots=a_{n}=0$, we define the operator $\theta:\Sigma^{n}\times\mathbb{Z}_{k}\to\Sigma^{n}\times\mathbb{Z}_{k}$,
$\theta\langle A,0\rangle=\langle[A_{i-1}(a_{i}-1)(s-1)^{j-i}]^{q}A_{n-qj},0\rangle,$
where,
if $n|k$ then $j=n$ and $q=1$. So, $\theta\langle A,0\rangle=\langle[A_{i-1}(a_{i}-1)(s-1)^{n-i},0\rangle$;
if $k|n$ and $n<k$ then $j$ is the smallest integer such that $k|j$ and $j\geq i$, and $q$ is the greatest integer such that $q\leq n/j$. Thus, in either case, $j=i+((n-i)\mod k)$.
The operator $\theta$ is applicable on any pair with second component 00 0, except for $\langle 0^{n},0\rangle$. For example, for $s=2$, $n=6$ and $k=2$, $\theta\langle 010000,0\rangle=\langle 000000,0\rangle$; $\theta\langle 011000,0\rangle=\langle 010101,0\rangle$; $\theta\langle 011101,0\rangle=\langle 011100,0\rangle$.
Definition (value $T$).
Let $T$ be the integer such that $\theta^{T}\langle(s-1)^{n},0\rangle=\langle(0^{n},0\rangle)$.
Lemma 2.
The list $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$ is strictly decreasing in $\succ$.
Proof.
Let’s see that for every pair ${\mathcal{A}}=\langle A,0\rangle$, ${\mathcal{A}}\succ\theta{\mathcal{A}}$. Using the definition of $\theta$,
$\langle A,0\rangle\succ\langle[A_{i-1}(a_{i}-1)(s-1)^{j-i}]^{q}A_{n-qj},0\rangle.$
Since both pairs have second component 00 0, there is some $i$ such that
${a}_{1}\cdots{a}_{i}=A_{i-1}a_{i}>A_{i-1}(a_{i}-1).$
We conclude that $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$ is strictly decreasing in $\succ$. ∎
Observation 6.
When $n|k$ the operator $\theta$ yields a bijection between maximal pairs.
Proof.
Every pair of the form $\langle A,0\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ is maximal because, when $n$ divides $k$, there is just this unique rotation with second component 00 0. The definition of $\theta$ ensures that $\theta$ goes through all the pairs of the form $\langle A,0\rangle$ in $\succ$-decreasing order. The operator $\theta$ can be used forward for every pair $\langle A,0\rangle$ except for $\langle 0^{n},0\rangle$, and it can be used backwards for every pair $\langle A,0\rangle$ except for $\langle 0^{n},0\rangle$. Thus, except for the extremes, we can obtain the successor and the predecessor of a maximal pair in the order $\succ$. ∎
When $k|n$ and $k<n$ the operator $\theta$ is not injective nor surjective over pairs with second component 00 0. For example, for $s=2$, $n=4$ and $k=2$, we see $\theta$ is not injective because $\theta\langle 0100,0\rangle=\langle 0000,0\rangle$ and also $\theta\langle 0001,0\rangle=\langle 0000,0\rangle$. To see that $\theta$ is not surjective observe that the pair ${\mathcal{B}}=\langle 1011,0\rangle$ is not in the image of $\theta$, because there is no pair ${\mathcal{A}}$ such that $\theta{\mathcal{A}}={\mathcal{B}}$.
It is possible to construct the reverse of list $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$. In case $n|k$ the operator $\theta$ defines a bijection between maximal pairs. In case $k|n$ and $k<n$, $\theta$ is not injective, there are pairs that have more than one preimage by $\theta$. However, except for $\langle(s-1)^{n},0\rangle$ every element has one predecessor in the list $(\theta^{i}(\langle s-1)^{n},0\rangle)_{0\leq i\leq T}$, which is just one of the possible preimages by $\theta$.
Lemma 3.
Let $n$ and $k$ positive integers such that $k|n$. Every element ${\mathcal{A}}=\langle A,0\rangle$ in $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$, except $\langle(s-1)^{n},0\rangle$, has a predecessor which is the preimage of ${\mathcal{A}}$ by $\theta$ given by
$\langle A_{u-1}(a_{u}+1))0^{n-u},0\rangle$
where
$A=\left(A_{u}(s-1)^{r-u}\right)^{w}A_{v}$
$a_{u}<(s-1)$, and $r$ is the smallest multiple of $k$ such that $v<r$ and $r-k\leq u\leq r$.
Proof.
First notice that this factorization always exists
$A=(A_{r})^{w}A_{v}=(A_{u}(s-1)^{r-u})^{w}A_{v}.$
If $w=1$ the $r=n$, $v=0$ and $A=A_{n}=A_{r}=A_{u}(s-1)^{r-u}$. Let ${\mathcal{B}}=\langle B,0\rangle$ be the pair obtained by undoing the transformation done by the $\theta$ operator, knowing that ${\mathcal{A}}$ and ${\mathcal{B}}$ are in $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$,
$\theta{\mathcal{B}}=\langle[B_{i-1}({b}_{i}-1)(s-1)^{j-i}]^{q}B_{n-qj},0\rangle={\mathcal{A}}$
where $i$ satisfies $b_{i+1}=\ldots b_{n}=0$, and $i\leq j$ and $k|j$. Thus,
$[B_{i-1}(b_{i}-1)(s-1)^{j-i}]^{q}B_{n-qj}=(A_{u}(s-1)^{r-u})^{w}A_{v}.$
The word $B$ is determined by $r=j,u=i,w=q,v=n-qj$ and $A_{i-1}(a_{i}+1)=B_{i}$. Then,
$\displaystyle\langle B_{i-1}b_{i}0^{n-j},0\rangle=\langle A_{u-1}(a_{u}+1))0^{n-u},0\rangle,$
$\displaystyle\langle[B_{i-1}({b}_{i}-1)(s-1)^{j-i}]^{q}B_{n-qj},0\rangle=[A_{u-1}a_{u}(s-1)^{r-u}]^{w}A_{v}.$
Lemma 4.
If ${\mathcal{A}}\succ{\mathcal{B}}\succ{\mathcal{C}}$, with ${\mathcal{C}}=\theta{\mathcal{A}}$, then ${\mathcal{B}}$ is not maximal .
Proof.
Let ${\mathcal{A}}=\langle A,0\rangle$, ${\mathcal{B}}=\langle B,0\rangle$, ${\mathcal{C}}=\theta{\mathcal{A}}=\langle C,0\rangle$ and indices $i,j$ such that
$\displaystyle A$ $\displaystyle$== $\displaystyle{a}_{1}\cdots{a}_{i}0^{n-i},\text{with $\displaystyle a_{i}>0$},$
$\displaystyle B$ $\displaystyle$== $\displaystyle{b}_{1}\cdots{b}_{n},$
$\displaystyle C$ $\displaystyle$== $\displaystyle{c}_{1}\cdots{c}_{n}=[{a}_{1}\cdots{a}_{i-1}(a_{i}-1)(s-1)^{j-i}]^{q}{a}_{1}\cdots{a}_{n-qj}.$
Assume ${\mathcal{A}}\succ{\mathcal{B}}\succ{\mathcal{C}}$ and, by way of contradiction suppose ${\mathcal{B}}$ is a maximal pair. Since ${\mathcal{A}}\succ{\mathcal{B}}$,
$b_{1}=a_{1},\quad b_{2}=a_{2},\quad\ldots,\quad b_{i-1}=a_{i-1},$
and it should be $a_{i}>b_{i}$. Since ${\mathcal{B}}\succ{\mathcal{C}}$ then $b_{i}\geq c_{i}=a_{i}-1$; hence, $b_{i}=a_{i}-1$. Then, we have $B_{i}=C_{i}$. And for $\ell=1,\ldots j-i$ we have $b_{i+\ell}=s-1$. This is because ${\mathcal{B}}\succ{\mathcal{C}}$ and ${c}_{i+1}\cdots{c}_{i+\ell}=(s-1)^{j-i}$, then ${b}_{i+1}\cdots{b}_{i+\ell}=(s-1)^{j-i}$, given that $s-1$ is the lexicographically greatest symbol.
We now show that the other symbols in ${\mathcal{B}}$ and ${\mathcal{C}}$ also coincide. Since ${\mathcal{B}}$ is a maximal pair, $B\geq{b}_{j+1}\cdots{b}_{n}{b}_{1}\cdots{b}_{j}$ and we know that $a_{1}=b_{1}\geq b_{j+1}$. Since ${\mathcal{B}}\succ{\mathcal{C}}$ we have $b_{j+1}\geq c_{j+1}=a_{1}$, hence $b_{j+1}=a_{1}$. Repeating this argument we obtain for $m=1,\ldots,q$ and $p=1,\ldots,i-1$,
$b_{mj+p}=a_{p}$, for $mj+p\leq n$,
$b_{mj+i}=a_{i}-1$ and
$b_{mj+i+\ell}=s-1$ for $\ell=1,\ldots,j-i$.
This would imply ${\mathcal{B}}={\mathcal{C}}=\theta{\mathcal{A}}$. Thus, there is no maximal pair ${\mathcal{B}}$ such that ${\mathcal{A}}\succ{\mathcal{B}}\succ{\mathcal{C}}$. ∎
3. Statement of Theorem 1
Definition (list ${\mathcal{M}}$ of maximal pairs of $\Sigma^{n}\times\mathbb{Z}_{k}$).
Let $n$ and $k$ be positive integers such that $n|k$ or $k|n$. Define ${\mathcal{M}}$ by removing from $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$ the pairs that are not maximal. In case $n|k$ all the elements are maximal, none is removed. In case $k|n$ with $k<n$, for each pair ${\mathcal{A}}$ in $(\theta^{i}\langle(s-1)^{n},0\rangle)_{0\leq i\leq T}$, except ${\mathcal{A}}=\langle 0^{n},0\rangle$, remove $\theta{\mathcal{A}}$, $\theta^{2}{\mathcal{A}},\ldots\theta^{h-1}{\mathcal{A}}$ where $h$ is the least such that $\theta^{h}({\mathcal{A}})$ is maximal . Let $M$ be the number of elements of the list ${\mathcal{M}}$.
Definition (list ${\mathcal{L}}$ of Lyndon pairs).
Let $n$ and $k$ be positive integers such that $k|n$ or $n|k$. We define the list ${\mathcal{L}}$ as the list of Lyndon pairs of the elements in the list ${\mathcal{M}}$. Since ${\mathcal{M}}$ has $M$ elements, ${\mathcal{L}}$ has $M$ elements as well.
Recall that a necklace is a circular word and a necklace is $(n,k)$-perfect if all words of length $n$ occur $k$ times but at positions with different congruence modulo $k$, for any convention of the starting position.
Theorem 1.
Let $n$ and $k$ be positive integers such that $k|n$ or $n|k$. Let $X$ be the concatenation of all the words in the Lyndon pairs in the order given by ${\mathcal{L}}$. Then, $X$ is the lexicographically greatest $(n,k)$-perfect necklace.
Here is an example for $s=2$, $n=6$ and $k=2$. The lexicographically greatest $(n,k)$-perfect necklace is obtained by concatenating the following words (the symbol $\mid$ is for ease of reading):
11111110111101111100111010111001111000110110110101110100110010110001110000101010011010001001011001001000011000000101010001000000conditional11delimited-∣∣111110111101delimited-∣∣111100111010delimited-∣∣111001111000delimited-∣∣110110110101delimited-∣∣110100110010delimited-∣∣110001110000delimited-∣∣10101001delimited-∣∣101000100101delimited-∣∣100100100001delimited-∣∣10000001delimited-∣∣01010001000000 11$\mid$ 111110$\mid$ 111101$\mid$ 111100$\mid$ 111010$\mid$ 111001$\mid$ 11% 1000$\mid$ 110110$\mid$ 110101$\mid$ 110100$\mid$ 110010$\mid$ 110001$\mid$ 110000$\mid$ 10% $\mid$ 101001$\mid$ 101000$\mid$ 100101$\mid$ 100100$\mid$ 100001$\mid$ 100000$\mid$ 01$\mid$ 0% 10100$\mid$ 010000$\mid$ 0011 ∣ 111110 ∣ 111101 ∣ 111100 ∣ 111010 ∣ 111001 ∣ 111000 ∣ 110110 ∣ 110101 ∣ 110100 ∣ 110010 ∣ 110001 ∣ 110000 ∣ 10 ∣ 101001 ∣ 101000 ∣ 100101 ∣ 100100 ∣ 100001 ∣ 100000 ∣ 01 ∣ $0$ $0$ 00.
4. Proof of Theorem 1
We divide the proof of Theorem 1 in three parts:
in Section 4.1 we show that length of $X$ is $s^{n}k$,
in Section 4.2 we prove that $X$ is $(n,k)$-perfect, and
in Section 4.3 we show that $X$ is the lexicographically greatest $(n,k)$-perfect necklace.
We start with two lemmas about the list ${\mathcal{M}}$ of maximal pairs.
Lemma 5.
The list ${\mathcal{M}}$ starts with the $\succ$-greatest pair $\langle(s-1)^{n},0\rangle$, ends with the $\succ$-smallest pair $\langle 0^{n},0\rangle$, and contains all maximal pairs in strictly decreasing $\succ$-order.
Proof of Lemma 5.
Case $k\mid n,k<n$.
  • By definition, the list ${\mathcal{M}}$ starts with $\langle(s-1)^{n},0\rangle$.
  • The maximal pairs are in $\succ$-decreasing order: This is because the list ${\mathcal{M}}$ is constructed by successive applications of the operator $\theta$ and by Lemma 2 the list is strictly $\succ$-decreasing.
  • No maximal pair is missing: Suppose ${\mathcal{A}}$ is in ${\mathcal{M}}$ and let $h$ be the least such that $\theta^{h}({\mathcal{A}})$ is maximal . To argue by contradiction, suppose there is a maximal pair ${\mathcal{B}}$ such that ${\mathcal{A}}\succ{\mathcal{B}}\succ\theta^{h}({\mathcal{A}})$. Then, there is $i$, $0\leq i<h$, such that
    ${\mathcal{A}}\succ\theta{\mathcal{A}}\succ\ldots\succ\theta^{i}{{\mathcal{A}}}\succ{\mathcal{B}}\succ\theta^{i+1}{{\mathcal{A}}}\succ\ldots\succ\theta^{h-1}{{\mathcal{A}}}\succ\theta^{h}{{\mathcal{A}}}.$
    Thus, ${\mathcal{B}}$ occurs in between some ${\mathcal{D}}$ and $\theta{\mathcal{D}}$. By Lemma 4 this is impossible.
  • The list ${\mathcal{M}}$ ends with $\langle 0^{n},0\rangle$: There is no ${\mathcal{A}}$ such that $\langle 0^{n},0\rangle\succ{\mathcal{A}}$, and the list ${\mathcal{L}}$ is strictly $\succ$-decreasing. By Lemma 2, the operator $\theta$ applies to any pair except $\langle 0^{n},0\rangle$.
Case $n\mid k$, $n\leq k$. It is the same proof as in the previous case, but simpler because $\theta$ yields exactly all the maximal pairs in $\succ$-decreasing order. ∎
Example 1.
Let $s=2$, $n=6$ and $k=2$. The list ${\mathcal{M}}$ of maximal pairs is:
111111,0,111110,0,111101,0,111100,0,111010,0,111001,0,111000,0,110110,0,110101,0,110100,0,110010,0,110001,0,110000,0,101010,0,101001,0,101000,0,100101,0,100100,0,100001,0,100000,0,010101,0,010100,0,010000,0,000000,0111111011111001111010111100011101001110010111000011011001101010110100011001001100010110000010101001010010101000010010101001000100001010000000101010010100001000000000000$\langle$ 111111,0$\rangle$,$\langle$ 111110,0$\rangle$,$\langle$ 111101,0% $\rangle$,$\langle$ 111100,0$\rangle$,$\langle$ 111010,0$\rangle$,$\langle$ 111001,0% $\rangle$,$\langle$ 111000,0$\rangle$,$\langle$ 110110,0$\rangle$,\\ $\langle$ 110101,0$\rangle$,$\langle$ 110100,0$\rangle$,$\langle$ 110010,0$\rangle$,$\langle$ 110001,0$\rangle$,$\langle$ 110000,0$\rangle$,$\langle$ 101010,0$\rangle$,$\langle$ 101001,0$\rangle$,$\langle$ 101000,0$\rangle$,\\ $\langle$ 100101,0$\rangle$,$\langle$ 100100,0$\rangle$,$\langle$ 100001,0$\rangle$,$\langle$ 100000,0$\rangle$,$\langle$ 010101,0$\rangle$,$\langle$ 010100,0$\rangle$,$\langle$ 010000,0$\rangle$,$\langle$ 000000,0$\rangle$ 111111 , 0 ⟩ , ⟨ 111110 , 0 ⟩ , ⟨ 111101 , 0 ⟩ , ⟨ 111100 , 0 ⟩ , ⟨ 111010 , 0 ⟩ , ⟨ 111001 , 0 ⟩ , ⟨ 111000 , 0 ⟩ , ⟨ 110110 , 0 ⟩ , ⟨ 110101 , 0 ⟩ , ⟨ 110100 , 0 ⟩ , ⟨ 110010 , 0 ⟩ , ⟨ 110001 , 0 ⟩ , ⟨ 110000 , 0 ⟩ , ⟨ 101010 , 0 ⟩ , ⟨ 101001 , 0 ⟩ , ⟨ 101000 , 0 ⟩ , ⟨ 100101 , 0 ⟩ , ⟨ 100100 , 0 ⟩ , ⟨ 100001 , 0 ⟩ , ⟨ 100000 , 0 ⟩ , ⟨ 010101 , 0 ⟩ , ⟨ 010100 , 0 ⟩ , ⟨ 010000 , 0 ⟩ , ⟨ 000000 , 0 ⟩.
The next is the key lemma.
Lemma 6.
Assume $k|n$. If ${\mathcal{A}}=\langle A,0\rangle$ is followed by ${\mathcal{B}}=\langle B,0\rangle$ in the list of maximal pairs ${\mathcal{M}}$ then $A$ is a prefix of $\overline{A}\,\overline{B}$.
Proof.
We can write ${\mathcal{A}}$ as
$\overline{{\mathcal{A}}}^{q}=\langle(A_{i}0^{p-i})^{q},0\rangle,$
where $q=n/p$ and $a_{i}>0$. If $q=1$ then $\overline{{\mathcal{A}}}={\mathcal{A}}=\langle A,0\rangle$ and $\overline{{\mathcal{A}}}\,\overline{{\mathcal{B}}}=\langle A,0\rangle\langle\overline{B},0\rangle=\langle A\overline{B},0\rangle$. Otherwise, $q>1$ and, since ${\mathcal{B}}=\theta{\mathcal{A}}^{h}$ for the smallest $h$ such that it is maximal, the shape of ${\mathcal{B}}$ is
${\mathcal{B}}=\langle\overline{A}^{q-1}A_{i-1}(a_{i}-1)(s-1)^{j-i}C,0\rangle$
for some word $C$ and for $j$ the smallest such that $i\leq j\leq p$ and $k\mid j$. Since ${\mathcal{B}}=\langle B,0\rangle$ and $B$ starts with $\overline{A}^{q-1}A_{i-1}(a_{i}-1)$ we have $\langle\overline{B},0\rangle=\langle B,0\rangle$, hence $\overline{{\mathcal{B}}}={\mathcal{B}}$. Then,
$\displaystyle\overline{{\mathcal{A}}}\,\overline{{\mathcal{B}}}=\langle\overline{A},0\rangle\langle\overline{B},0\rangle=\langle\overline{A},0\rangle\langle B,0\rangle=\langle\overline{A}B,0\rangle=\langle\overline{A}\,\overline{A}^{q-1}A_{i-1}(a_{i}-1)(s-1)^{j-i}C,0\rangle.$
In both cases $A$ is prefix of $\overline{A}\,\overline{B}$. ∎
4.1. Proof that necklace $X$ has length $s^{n}k$
4.1.1. When $k|n,k<n$
The necklace $X$ is the concatenation of all the words of the reduced maximal pairs exactly once. Thus, the length of $X$ is
$|X|=\sum_{\overline{{\mathcal{A}}}\text{ is Lyndon }}|\overline{{\mathcal{A}}}|.$
Each $\overline{{\mathcal{A}}}$ is an element of ${\mathcal{L}}$, which consists of the reductions of the pairs in the list ${\mathcal{M}}$, which by Lemma 5 contains all the maximal pairs. The length of $\overline{A}=\langle\overline{A},0\rangle$ is the length of the word $\overline{A}$, which is the number of different rotations of ${\mathcal{A}}$, see Observation 5. If $\overline{{\mathcal{A}}}=\langle\overline{A},0\rangle=\langle{a}_{1}\cdots{a}_{p},0\rangle$, then ${\mathcal{A}}$ has $p$ different rotations: $\langle{a}_{1}\cdots{a}_{p},0\rangle$, $\langle{a}_{2}\cdots{a}_{n}a_{1},1\rangle$,…, $\langle a_{p}{a}_{1}\cdots{a}_{n-1},p-1\rangle$. So, each of the $p$ positions in $\overline{A}$ is the start of a different rotation of ${\mathcal{A}}$. By Lemma 6 if the successor of ${\mathcal{A}}=\langle A,0\rangle$ in ${\mathcal{M}}$ is ${\mathcal{B}}=\langle B,0\rangle$ then in $X$ we have $\overline{A}\,\overline{B}$, and $A$ is a prefix of $\overline{A}\,\overline{B}$. Let $X=a_{0}\ldots a_{\ell-1}$. We argue that each position $i$, for $i=0,1,\ldots,\ell-1$, is the start of one of the different pairs $\langle{a}_{i}\cdots{a}_{i+n-1},i\mod k\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$. We conclude that $X$ has exactly one position for each of the different pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$. Thus, the length of $X$ is $s^{n}k$.
4.1.2. When $n|k$, $n\leq k$
The necklace $X$ is the concatenation of all the words of the expanded maximal pairs. Thus, length of $X$ is
$|X|=\sum_{\widetilde{{\mathcal{A}}}\text{ is Lyndon }}|\widetilde{{\mathcal{A}}}|.$
The length of a Lyndon pair $\widetilde{A}=\langle\widetilde{A},0\rangle$ is the length of the expansion of a maximal pair $\langle A,0\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$, which is exactly $k$. There are $s^{n}$ many Lyndon pairs because, given that $n|k$, every $\langle A,0\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ is maximal. Since $X$ is the concatenation of all the Lyndon pairs, the length of $X$ is $s^{n}k$.
4.2. Proof that necklace $X$ is $(n,k)$-perfect
We need to show that each word of length $n$ occurs exactly $k$ times, at positions with different congruence modulo $k$. In this proof we number the positions of $X$ starting at 00 0; this is convenient for the presentation because the positions with congruence 00 0 are multiple of $k$. We say that we find a pair $\langle A,u\rangle$ in $X=a_{0}a_{1}\ldots a_{s^{n}k-1}$ when there is a position $p$ such that $0\leq p<s^{n}k$, $p\equiv u\pmod{k}$, and $a_{p}\ldots a_{p+|A|-1}=A$. To prove that $X$ is a $(n,k)$-perfect necklace we need to find all the rotations of all the maximal pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$ in the necklace $X$.
4.2.1. When $k|n,k<n$
Each maximal pair ${\mathcal{A}}=\langle A,0\rangle$ has exactly $|\overline{A}|$ many rotations.
Case $A=(s-1)^{n}$.
Since $k|n$, $\overline{{\mathcal{A}}}$=$\langle\overline{A},0\rangle=\langle(s-1)^{k},0\rangle$, application of $\theta$ on ${\mathcal{A}}$ yields the maximal pair
$\theta{\mathcal{A}}={\mathcal{B}}=\langle B,0\rangle=\langle(s-1)^{n-1}(s-2),0\rangle.$
Since $\overline{{\mathcal{A}}}$ and $\overline{{\mathcal{B}}}$ are consecutive Lyndon pairs in ${\mathcal{L}}$, the construction of $X$ puts $\overline{A}$ followed by $\overline{B}$,
$\overline{{\mathcal{A}}}\,\overline{{\mathcal{B}}}=\langle(s-1)^{k},0\rangle\langle(s-1)^{n-1}(s-2),0\rangle=\langle(s-1)^{n+k-1}(s-2),0\rangle.$
This yields all rotations of ${\mathcal{A}}$: $\langle(s-1)^{n},0\rangle$, …, $\langle(s-1)^{n},k-1\rangle$.
Case $A=0^{n}$.
Since $k|n$, $\overline{{\mathcal{A}}}=\langle\overline{A},0\rangle=\langle 0^{k},0\rangle$. Consider the maximal pairs for $i$=0,$1\displaystyle 0,1,\ldots,k-$ 1,
${\mathcal{B}}=\langle B,0\rangle=\langle 0^{i}10^{k-1-i}0^{n-k},0\rangle.$
By Lemma 1, for each of these ${\mathcal{B}}$ the next maximal pair in ${\mathcal{M}}$ is
$\theta{\mathcal{B}}={\mathcal{C}}=\langle C,0\rangle=\langle(0^{i+1}(s-1)^{k-1-i})^{n/k},0\rangle.$
Since $\overline{{\mathcal{C}}}=\langle 0^{i+1}(s-1)^{k-1-i},0\rangle$, it is clear that ${\mathcal{C}}$ is maximal because all the rotations of ${\mathcal{C}}$ that have second component 00 0 are identical to ${\mathcal{C}}$. Since $\overline{{\mathcal{B}}}$ and $\overline{{\mathcal{C}}}$ are consecutive Lyndon pairs in ${\mathcal{L}}$, the construction of $X$ puts $\overline{B}$ followed by $\overline{C}$. Notice that for each $i=0,1,\ldots,k-1$,
¯$\overline{{\mathcal{B}}}\,\overline{{\mathcal{C}}}=\langle 0^{i}10^{k-1-i}0^{n-k}0^{i+1}(s-1)^{k-i-1},0\rangle=\langle 0^{i}10^{n}(s-1)^{k-i-1},0\rangle$
gives rise to the pair $\langle 0^{n},i+1\rangle$. Thus, we have all the rotations ${\mathcal{A}}=\langle 0^{n},0\rangle$, which are $\langle 0^{n},0\rangle$, $\ldots$, $\langle 0^{n},k-1\rangle$.
Case $(s-1)^{n}>A>0^{n}$
Every maximal pair ${\mathcal{A}}=\langle A,0\rangle$ different from $\langle 0^{n},0\rangle$ has the form
${\mathcal{A}}=\langle(A_{p})^{q+1},0\rangle,$
where $A_{p}$ is $\overline{A}$ with $p$ the smallest integer such that $A=(A_{p})^{q+1}$ and $q+1=(n/p)$.
Subcase $q>0$. The pair ${\mathcal{A}}$, different from $\langle 0^{n},0\rangle$, always has a successor maximal pair in the list ${\mathcal{M}}$:
${\mathcal{B}}=\theta{\mathcal{A}}=\langle(A_{p})^{q}A_{i-1}(a_{i}-1)(s-1)^{j-i}{b}_{i+1}\cdots{b}_{p},0\rangle,$
where $i$ is the largest with $a_{i}>0$, $j$ is the smallest multiple of $k$ with $j\geq i$. Notice that $A_{p}\neq 0^{p}$. Since $\overline{{\mathcal{B}}}={\mathcal{B}}$,
$\overline{{\mathcal{A}}}\,\overline{{\mathcal{B}}}=\langle(A_{p})^{q+1}A_{i-1}(a_{i}-1)(s-1)^{j-i}{b}_{i+1}\cdots{b}_{p},0\rangle,$
and it also yields the first $i$ left rotations of ${\mathcal{A}}$, which are $\langle{a}_{r+1}\cdots{a}_{n}A_{r},r\rangle$, with $0\leq r<i$. It remains to identify in the constructed necklace $X$ the $p-i$ right rotations of ${\mathcal{A}}=\langle(A_{p})^{q+1},0\rangle$. These are of the form
$\langle a_{r}\ldots a_{p}A_{p}^{q}A_{p-i-r},-r\rangle,$
for $1\leq r\leq p-i$. By Lemma 3 we can consider the predecessor ${\mathcal{Q}}$ of ${\mathcal{A}}$ by $\theta$,
$\theta{\mathcal{Q}}={\mathcal{A}}=\langle(A_{p})^{q+1},0\rangle$
where
${\mathcal{Q}}=\langle A_{p-1}(a_{p}+1)0^{pq},0\rangle.$
To see that ${\mathcal{Q}}$ is a maximal pair consider first ${\mathcal{R}}=\langle A_{p},0\rangle$ which, by Observation 2, is maximal. We now argue that $A_{p}$ has no proper suffix ${a}_{i}\cdots{a}_{p}$ that coincides with $A_{p-i+1}$. If there were such a prefix we could construct the rotation of ${\mathcal{R}}$ given by the pair ${\mathcal{S}}=\langle{a}_{i}\cdots{a}_{p}A_{i-1},0\rangle$ and one of the following would be true:
  • ${\mathcal{S}}={\mathcal{R}}$: but this is impossible by Observation 5.
  • ${\mathcal{R}}\succ{\mathcal{S}}$: Since ${\mathcal{R}}=\langle A_{p-i+1}{a}_{p-i+2}\cdots{a}_{p},0\rangle$, ${\mathcal{S}}=\langle a_{i}\ldots a_{p}A_{i-1},0\rangle$ and we assumed ${a}_{i}\cdots{a}_{p}=A_{p-i+1}$, necessarily ${a}_{p-i+2}\cdots{a}_{p}>A_{i-1}$. But this contradicts that ${\mathcal{R}}$ is a maximal pair.
  • ${\mathcal{S}}\succ{\mathcal{R}}$: There is a rotation of ${\mathcal{R}}$ which is $\succ$-greater than ${\mathcal{R}}$, contradicting that ${\mathcal{R}}$ is a maximal pair.
We conclude that all suffixes ${a}_{i}\cdots{a}_{p}$ of $R$ are lexicographically smaller than $A_{p-i}$. Therefore, all suffixes of ${a}_{i}\cdots{a}_{p-1}(a_{p}+1)$ of $Q$ are lexicographically smaller than or equal to $A_{p-i}$. We already argued that ${\mathcal{Q}}$ is indeed maximal. For any rotation of ${\mathcal{Q}}$ of the form $\langle{a}_{i}\cdots{a}_{p-1}(a_{p}+1)0^{pq}{a}_{1}\cdots{a}_{i-1},0\rangle$, for $i>1$, we argued that ${a}_{i}\cdots{a}_{p-1}(a_{p}+1)$ is lexicographically smaller than or equal to $A_{p-i}$. Moreover, ${a}_{i}\cdots{a}_{p-1}(a_{p}+1)0^{pq}$ is lexicographically smaller than $Q=A_{p-1}(a_{p}+1)0^{pq}$. For any rotation of ${\mathcal{Q}}$ that starts with a suffix of $0^{pq}$ it can not be maximal, because for any $m$, $Q_{m}>0^{m}$, otherwise ${\mathcal{A}}$ would not be maximal.
We have the maximal successive pairs ${\mathcal{Q}}=\langle Q,0\rangle$, $\theta{\mathcal{Q}}={\mathcal{A}}=\langle A,0\rangle$ and ${\mathcal{B}}=\theta{\mathcal{A}}=\langle B,0\rangle$. From the arguments above, $\overline{Q}=Q$ and $\overline{B}=B$. Since ${\mathcal{A}}=\langle(A_{p})^{q+1},0\rangle$ and we assumed $q>1$, $A$ is not equal to $\overline{A}$. Then, $\overline{Q}\,\overline{A}\,\overline{B}=Q\overline{A}B$. Observe that the last $pq$ symbols of $Q$ followed by $\overline{A}$, followed by the first $pq$ symbols of $B$ give rise to
$\langle 0^{pq}(A_{p})^{q+1},0\rangle,$
where 00 0 is because $p$ is a multiple of $k$. We now identify in this pair the $p-i$ rotations of ${\mathcal{A}}$ to the right. Since ${\mathcal{A}}=\langle A,0\rangle$ where $A=(A_{i}0^{p-i})^{q+1}$, after $r$ rotations to the right of ${\mathcal{A}}$, with $1\leq r\leq p-i$, we obtain
$\langle 0^{r}(A_{p})^{q}A_{i}0^{p-i-r},-r\rangle.$
Subcase $q=0$. We need to see that for the maximal pairs ${\mathcal{A}}=\langle A,0\rangle$ where $A$ is reduced, that is $A=\overline{A}$, we can find all rotations of ${\mathcal{A}}$ in the constructed $X$. If $A=0^{k-1}10^{n-k}$ then $\theta{\mathcal{A}}={\mathcal{B}}=\langle B,0\rangle=\langle 0^{n},0\rangle$, which is the last maximal pair in the list ${\mathcal{M}}$ and $\overline{B}=0^{k}$. Then,
$\overline{A}\,\overline{B}=0^{k-1}10^{n-k}0^{k}=0^{k-1}10^{n},$
and we can find $k$ left rotations of ${\mathcal{A}}$ which are of the form
$\langle 0^{k-1-r}10^{n-k+r},r\rangle\text{ for $\displaystyle r=0,1,\ldots k-1$}.$
Now assume $A\neq 0^{k-1}10^{n-k}$. Let’s write
${\mathcal{A}}=\langle A_{i}0^{n-i},0\rangle,$
with $i$ such that $a_{i}>a_{i+1}=\ldots=a_{n}=0$, ${\mathcal{B}}=\langle B,0\rangle=\theta{\mathcal{A}}$ and ${\mathcal{C}}=\langle C,0\rangle=\theta{\mathcal{B}}=\theta^{2}{{\mathcal{A}}}$. Then ${\mathcal{B}}$ is of the form
${\mathcal{B}}=\langle A_{i-1}(a_{i}-1)(s-1)^{j-i}{b}_{j+1}\cdots{b}_{p},0\rangle,$
with $j$ is the least multiple of $k$ with $j\geq i$. Since $A=\overline{A}$, and by Lemma 6 $B$ is a prefix of $\overline{B}\,\overline{C}$, then $AB$ is a prefix of $\overline{A}\,\overline{B}\,\overline{C}$, and we can find the first $i$ left rotations of ${\mathcal{A}}$ in it, which are of the form
$\langle{a}_{r+1}\cdots{a}_{n}A_{r},r\rangle,\text{ for $\displaystyle r=0,1,\ldots,i-1$}.$
It remains to find $n-i$ right rotations of ${\mathcal{A}}$, which are of the form
$\langle 0^{r}A_{i}0^{n-i-r},-r\rangle$
for $r=1,\ldots,n-i$. Equivalently we can write it as
$\langle 0^{n-i-h}A_{i}0^{h},(n-i-h)\rangle$
for $h=0,1,\ldots,n-i-1$.
If $h=0$ and $A_{i}=(s-1)^{i}$ then the pair is $\langle 0^{n-i}(s-1)^{i},n-i\rangle$. We find it in the pair $\langle 0^{n}(s-1)^{n},0\rangle$, which results from the concatenation of the words in the last two Lyndon pairs in ${\mathcal{L}}$ and the first two, that is ${\mathcal{L}}_{M-1}$ and ${\mathcal{L}}_{M}$, followed by ${\mathcal{L}}_{1}$ and ${\mathcal{L}}_{2}$.
If $h\neq 0$ or $A_{i}\neq(s-1)^{i}$ then we find the right rotations of ${\mathcal{A}}$ in the concatenation of the words of three Lyndon pairs, that we call $\overline{{\mathcal{Q}}_{t}}\,\overline{{\mathcal{P}}}\,\overline{{\mathcal{R}}}$ and we define below. Recall that ${\mathcal{A}}=\langle A,0\rangle=\langle A_{i}0^{n-i},0\rangle$, with $i$ such that $a_{i}>a_{i+1}=\ldots=a_{n}=0$, is a maximal pair. We claim there is a unique pair ${\mathcal{Q}}$ of the form
${\mathcal{Q}}=\langle A_{i}0^{h}S,0\rangle,$
where $|S|=n-(i+h)$ is not zero. Notice that ${\mathcal{Q}}$ may not be maximal.
Since ${\mathcal{Q}}$ is a pair $\succ$-greater than ${\mathcal{A}}$, we know ${\mathcal{Q}}$ is a predecessor of ${\mathcal{A}}$ by $\theta$. The pair ${\mathcal{Q}}_{t}$ is the closest predecessor of ${\mathcal{Q}}$ by $\theta$ that is a maximal pair. For this, consider the successive application of the operator $\theta$, which allows us to traverse the list $(\theta^{l}\langle(s-1)^{n},0\rangle)_{1\leq l\leq T}$. Our interest is to traverse it backwards. Here is a diagram:
${\mathcal{Q}}_{t}$${\mathcal{Q}}\!=\!\langle A_{i}0^{h}S,0\rangle$${\mathcal{A}}$apply many times $\theta$apply many times $\theta$
For $l=1,2,3,\ldots$ let ${\mathcal{Q}}_{l}=\langle Q_{l},0\rangle$ be the pair such that $\theta^{l}{{\mathcal{Q}}_{l}}={\mathcal{Q}}$. For $l=1$,
$\displaystyle{\mathcal{Q}}_{1}$ $\displaystyle=\langle Q_{1},0\rangle=\langle A_{u_{1}-1}(a_{u_{1}}+1)0^{n-u_{1}},0\rangle,$
$\displaystyle\theta{\mathcal{Q}}_{1}$ $\displaystyle=\langle(A_{u_{1}}(s-1)^{r_{1}-u_{1}})^{w_{1}}A_{v_{1}}S,0\rangle=\langle(A_{r_{1}})^{w_{1}}A_{v_{1}}S,0\rangle=\langle A_{i}0^{h}S,0\rangle={\mathcal{Q}},$
where $r_{1}$ is the least multiple of $k$ such that $v_{1}<r_{1}$, $a_{u_{1}}<(s-1)$, $A_{r_{1}}=A_{u_{1}}(s-1)^{r_{1}-u_{1}}$ and $r_{1}-k\leq u_{1}\leq r_{1}$. If ${\mathcal{Q}}_{1}$ is a maximal pair then fix $t=1$ and we have finished the search. Otherwise we consider the predecessor of ${\mathcal{Q}}_{1}$ by $\theta$,
${\mathcal{Q}}_{2}=\langle Q_{2},0\rangle=\langle A_{u_{2}-1}(a_{u_{2}}+1)0^{n-u_{2}},0\rangle$
such that
$\theta{\mathcal{Q}}_{2}=\langle(A_{u_{2}}(s-1)^{r_{2}-u_{2}})^{w_{2}}A_{v_{2}},0\rangle=\langle(A_{r_{2}})^{w_{2}}A_{v_{2}},0\rangle={\mathcal{Q}}_{1},$
where $r_{2}$ is the least multiple of $k$, $v_{2}<r_{2}$, and $a_{u_{2}}<(s-1)$ such that $A_{r_{2}}=A_{u_{2}}(s-1)^{r_{2}-u_{2}}$. We know that $a_{u_{2}}<(s-1)$ exists because ${\mathcal{Q}}_{1}\neq\langle(s-1)^{n},0\rangle$, otherwise ${\mathcal{Q}}_{1}$ would be a maximal pair, and then there is $a_{i}<s-1$ for some $1\leq i\leq r_{2}$. Notice that $u_{2}<u_{1}$. If ${\mathcal{Q}}_{2}$ is a maximal pair then we fix $t=2$ and we have finished the search. Otherwise, we repeat this procedure. In this way the predecessor of ${\mathcal{Q}}_{l-1}$ by $\theta$ is
${\mathcal{Q}}_{l}=\langle Q_{l},0\rangle=\langle A_{u_{l}-1}(a_{u_{l}}+1)0^{n-u_{l}},0\rangle,$
such that
$\theta{\mathcal{Q}}_{l}=\langle(A_{u_{l}}(s-1)^{r_{l}-u_{l}})^{w_{l}}A_{v_{l}},0\rangle=\langle(A_{r_{l}})^{w_{l}}A_{v_{l}},0\rangle={\mathcal{Q}}_{l-1},$
where $r_{l}$ is multiple of $k$, $v_{l}<r_{l}$, $u_{l}<u_{l-1}$ and $a_{u_{l}}<(s-1)$ such that
$A_{r_{l}}=A_{u_{l}}(s-1)^{r_{l}-u_{l}}.$
Eventually we find $t$ such that ${\mathcal{Q}}_{t}$ is a maximal pair. Consider now the three consecutive maximal pairs in the list ${\mathcal{M}}$,
${\mathcal{Q}}_{t}=\langle Q_{t},0\rangle,{\mathcal{P}}=\langle P,0\rangle,\text{ and }{\mathcal{R}}=\langle R,0\rangle.$
Let $\overline{{\mathcal{Q}}_{t}},\overline{{\mathcal{P}}}$ and $\overline{{\mathcal{R}}}$ be the corresponding Lyndon pairs. ${\mathcal{P}}$ always exists, because it’s either ${\mathcal{A}}$ or a maximal pair before it, and ${\mathcal{R}}$ is $\langle 0^{n},0\rangle$ in the worst case. Notice that $\overline{Q_{t}}$ ends with $0^{n-u_{t}}$ and, by Lemma 6, $\overline{P}\,\overline{R}$ starts with $P$. Therefore, $\overline{{\mathcal{Q}}_{t}}\,\overline{{\mathcal{P}}}\,\overline{{\mathcal{R}}}$ contains $0^{n-u_{t}}P=0^{n-u_{t}}A_{i}0^{h}C$, for some $C$. Since ${\mathcal{P}}$ is a maximal pair $\succ$-greater than or equal to ${\mathcal{A}}$, we can assert that its prefix is $A_{i}0^{h}$. Finally, we have
$u_{t}\leq u_{1}\leq i+h,$
because $u_{t}<u_{t-1}<\ldots<u_{1}$ and $u_{1}\leq i+h$ because $u_{1}$ was the position of a symbol in $A_{i}0^{h}$, which has length $i+h$. Then,
$\displaystyle u_{t}\leq u_{1}\leq i+h$ $\displaystyle\iff n-u_{t}\geq n-i-h.$
We conclude that $\overline{{\mathcal{Q}}_{t}}\overline{{\mathcal{P}}}\overline{{\mathcal{R}}}$ contains
$\langle 0^{n-i-h}A_{i}0^{h},(|\overline{Q_{h}}|-(n-i-h))\rangle.$
This can be rewritten as the pair
$\langle 0^{n-i-h}A_{i}0^{h},-(n-i-h)\rangle,$
because $|\overline{Q_{t}}|$ is a multiple of $k$, hence ${\mathcal{Q}}_{t}$ has second component 00 0.
4.2.2. When $n|k$, $n\leq k$
This proof is similar to the previous one, but now we need to use the corresponding definitions of $\theta$, the list ${\mathcal{M}}$ and the notion of expansion to define the list ${\mathcal{L}}$. The proof becomes simpler because it requires fewer cases, and each of these cases is simpler too. The key observation for this proof is that each word $A$ of length $n$, repeated $k/n$ times, determines a Lyndon pair $\widetilde{A}=\langle\widetilde{A},0\rangle$, which has exactly $|\widetilde{A}|=k$ many different rotations.
Case $A=0^{n}$.
We can find the first $k-n+1$ left rotations of $A$ inside $\widetilde{{\mathcal{A}}}=\langle\widetilde{A},0\rangle=\langle 0^{k},0\rangle$, which are of the form $\langle 0^{n},i\rangle$, with $1\displaystyle 0\leq i<k-n+$ 1. Consider the maximal pairs for $1\displaystyle 0\leq i<n-$ 1,
${\mathcal{B}}=\langle B,0\rangle=\langle 0^{i}10^{n-1-i},0\rangle.$
For each of these ${\mathcal{B}}$,
$\theta{\mathcal{B}}={\mathcal{C}}=\langle C,0\rangle=\langle 0^{i+1}(s-1)^{n-1-i},0\rangle.$
Since $\widetilde{{\mathcal{B}}}$ and $\widetilde{{\mathcal{C}}}$ are consecutive Lyndon pairs in ${\mathcal{L}}$, the construction of $X$ puts $\widetilde{B}$ followed by $\widetilde{C}$. Notice that for each $i=0,1,\ldots,k-1$,
~$\widetilde{{\mathcal{B}}}\,\widetilde{{\mathcal{C}}}=\langle(0^{i}10^{n-1-i})^{k/n-1}0^{i}10^{n}(s-1)^{n-1-i}(0^{i+1}(s-1)^{n-1-i})^{k/n-1},0\rangle$
gives rise to the pair $\langle 0^{n},i+1\rangle$. Thus, we have all the rotations of the pair ${\mathcal{A}}=\langle 0^{n},0\rangle$, which are $\langle 0^{n},0\rangle$, $\ldots$, $\langle 0^{n},k-1\rangle$.
Case $A>0^{n}$.
Let $A={a}_{1}\cdots{a}_{n}$, and let $j$ be the largest such that $a_{j}>0$. Fix $j$. Since the pair ${\mathcal{A}}$ is different from $\langle 0^{n},0\rangle$, it has necessarily a maximal pair successor in the list ${\mathcal{M}}$,
${\mathcal{B}}=\theta{\mathcal{A}}=\langle A_{j-1}(a_{j}-1)(s-1)^{n-j},0\rangle,$
Since ~$\widetilde{{\mathcal{B}}}={\mathcal{B}}^{k/n}$,
$\widetilde{{\mathcal{A}}}\,\widetilde{{\mathcal{B}}}=\langle A^{k/n}(A_{j-1}(a_{j}-1)(s-1)^{n-j})^{k/n},0\rangle,$
and it also yields the first $k-n+j$ left rotations of ${\mathcal{A}}$, which are $\langle{a}_{(i\mod n)+1}\cdots{a}_{n}A_{i\mod n},i\rangle$, for $0\leq i<k-n+j$. Note that all of the rotations for $A=(s-1)^{n}$ are contained inside these left rotations, as $j=n$, and then $0\leq i<k$.
If $A=(s-1)^{n}$ then the $k$ rotations of $\widetilde{{\mathcal{A}}}=\langle\widetilde{A},0\rangle$ are considered above by taking $j=n$ and $0\leq i<k$. In case $A\not=(s-1)^{n}$ we have considered just $k-n-j$ rotations which are
$\langle 0^{n-j-h}A_{j}0^{h},-(n-j-h)\rangle$
for $h=0,1,\ldots,n-j-1$.
If $h=0$ and $A_{j}=(s-1)^{j}$ then the pair is $\langle 0^{n-j}(s-1)^{j},j-n\rangle$. We find it in the pair $\langle 0^{k}(s-1)^{k},0\rangle$, which results from the concatenation of the last and first Lyndon pairs in ${\mathcal{L}}$, which are ${\mathcal{L}}_{M-1}$ and ${\mathcal{L}}_{1}$ respectively.
If $h\neq 0$ or $A_{j}\neq(s-1)^{j}$ then the pairs are $\langle 0^{n-j-h}A_{j}0^{h},-(n-j-h)\rangle$ for $h=1,\ldots n-j-1$, and we find them in the concatenation of the words of two Lyndon pairs, that we call $\widetilde{{\mathcal{Q}}}\,\widetilde{{\mathcal{P}}}$ and we define below. Let $i$ be the largest such that $a_{i}<s-1$ and $i\leq j+h$, and let ${\mathcal{Q}}=\langle Q,0\rangle=\langle A_{i-1}(a_{i}+1)0^{n-i},0\rangle$. By Observation 1, ${\mathcal{Q}}$ is a Lyndon pair, different from $\langle 0^{n},0\rangle$. Notice that ${\mathcal{Q}}$ is well defined because there is some $i$, with $1\leq i\leq n$ such that $1\displaystyle 1\leq a_{i}+1\leq s-$ 1. To see this notice that either $h=0$ but $A_{j}\neq(s-1)^{j}$, hence, there exists $a_{i}<(s-1)$; or $h>0$ and then $i=j+h$. Let ${\mathcal{P}}$ be the successor of ${\mathcal{Q}}$ in ${\mathcal{M}}$, which is of the form
$\langle A_{i}(s-1)^{n-i},0\rangle.$
The construction concatenates $\widetilde{{\mathcal{Q}}}\,\widetilde{{\mathcal{P}}}$, resulting in
$\langle A_{i-1}(a_{i}+1)0^{n-i},0\rangle^{k/n}\langle A_{i}(s-1)^{n-i},0\rangle^{k/n}.$
Finally, if we take the suffix $\widetilde{{\mathcal{Q}}}_{n-j-h}$ followed by the prefix $\widetilde{{\mathcal{P}}}_{n-j-h}$ it results in the pair $\langle 0^{n-j-h}A_{j+h},-(n-j-h)\rangle$. Given that we can do this for every possible $h=0,1,\ldots n-j-1$, we have obtained the wanted $n-j$ right rotations. Consequently, we have found all the rotations of ${\mathcal{A}}$.
4.3. Proof that necklace $X$ is the lexicographically greatest $(n,k)$-perfect necklace
The necklace $X$ is the concatenation of all the words in the Lyndon pairs in the list ${\mathcal{L}}=(\langle w_{i},0\rangle)_{1\leq i\leq M}$. Let $\ell_{0}=0$ and let $(\ell_{i})_{1\leq i\leq M}$ such that $\ell_{i}=\sum_{m=1}^{i}|w_{m}|$. We already showed $|X|=s^{n}k$, hence, $\ell_{M}=s^{n}k$. The necklace $X=w_{1}w_{2}\ldots w_{M}=a_{0}a_{1},\ldots a_{s^{n}k-1}$, where for $i=1,2,\ldots,M$, $w_{i}=a_{\ell_{i-1}},\ldots a_{\ell_{i}-1}$. Since each $w_{i}$ is a Lyndon word, $w_{i}$ is lexicographically greater than all of its rotations. Now, instead of looking at an index $i$ from 11 11 to $M$, let’s consider an index $j$ running through all the positions of $X$, from 00 0 to $s^{n}k-1$. Consider the set of pairs
${\mathcal{S}}=\{\langle a_{j}\ldots a_{j+n-1},j\mod k\rangle:0\leq j\leq s^{n}k-1\}.$
Since we already proved that $X$ is $(n,k)$-perfect, the set ${\mathcal{S}}$ is exactly $\Sigma^{n}\times\mathbb{Z}_{k}$. The construction guarantees that each position $j=0,\ldots,s^{n}k-1$ is the start of one of the different pairs in ${\mathcal{S}}$. As a consequence of the order of the elements in the list ${\mathcal{L}}$, when $j=\ell_{i}-1$ for some $i$, $1\leq i\leq|M|$, the pair $\langle a_{j}\ldots a_{j+n-1},0\rangle$ is the $\succ$-greatest among the remaining pairs with second component 00 0. If $k|n$ and $k<n$, each $a_{j}$ is a symbol in $n$ different pairs in $\Sigma^{n}\times\mathbb{Z}_{k}$, $n/k$ of them with second component 00 0. For example, $a_{j+n}$ is in $n$ pairs,
$\langle a_{j+1}\ldots a_{j+n},(j+1)\mod k\rangle,\ldots,\langle a_{j+n}\ldots a_{j+2n-1},(j+n)\mod k\rangle.$
If $n|k$ and $n\leq k$, each $a_{j}$ is a symbol in $k$ different pairs in $\Sigma^{k}\times\mathbb{Z}_{k}$. There is exactly one of these with second component 00 0 and it has the form $\langle A^{k/n},0\rangle$, with $A\in\Sigma^{n}$. For example, $a_{j+k}$ is in $k$ pairs,
$\langle a_{j+1}\ldots a_{j+k},(j+1)\mod k\rangle,\ldots,\langle a_{j+k}\ldots a_{j+2k-1},(j+k)\mod k\rangle.$
Implicitly our construction defines a function $f:\Sigma^{n}\to\Sigma$, such that for each $j=0,\ldots,s^{n}k-1$,
$a_{j+n}=f(a_{j},\ldots,a_{j+n-1}).$
Suppose $b_{0},b_{1},\ldots,b_{s^{n}k-1}$ is another $(n,k)$-perfect necklace and there is a position $p$ such that $b_{0}=a_{0},b_{1}=a_{1},\ldots,b_{p-1}=a_{p-1}$ but $b_{p}>a_{p}$. Then, there is a position $q$ congruent to 00 0 modulo $k$ such that $p-\max(n,k)+1\leq q\leq p+\max(n,k)-1$, and
${b}_{q}\cdots{b}_{q+\max(n,k)-1}>{a}_{q}\cdots{a}_{q+\max(n,k)-1}.$
Then, the respective pairs starting at position $q$ satisfy,
$\langle{b}_{q}\cdots{b}_{q+n-1},0\rangle\succ\langle{a}_{q}\cdots{a}_{q+n-1},0\rangle.$
But this contradicts that our construction puts in $a_{0}\ldots a_{s^{n}k-1}$ the pairs $\langle A,0\rangle$ in $\Sigma^{n}\times\mathbb{Z}_{k}$ in $\succ$-decreasing ordering. This completes the proof of Theorem 1.
Acknowledgements
We thank the anonymous reviewer for their careful reading of the manuscript and for the comments and suggestions, which significantly improved the presentation. This work is supported by the projects CONICET PIP 11220210100220CO and UBACyT 20020220100065BA.